### v1

# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

Abstract
We analyze generalizations of algorithms based on the short-path framework first proposed by Hastings [Quantum 2, 78 (2018)], which has been extended and shown by Dalzell et al. [STOC ’22] to achieve super-Grover speedups for certain binary optimization problems. We demonstrate that, under some commonly satisfied technical conditions, an appropriate generalization can achieve super-quadratic speedups not only over unstructured search but also over a classical optimization algorithm that searches for the optimum by drawing samples from the stationary distribution of a Markov Chain. We employ this framework to obtain algorithms for problems including variants of Max-Bisection, Max Independent Set, the Ising Model, and the Sherrington Kirkpatrick Model, whose runtimes are asymptotically faster than those obtainable from previous short path techniques. For random regular graphs of sufficiently high degree, our algorithm is super-quadratically faster than the best rigorously proven classical runtimes for regular graphs. Our results also shed light on the quantum nature of short path algorithms, by identifying a setting where our algorithm is super-quadratically faster than any polynomial time Gibbs sampler, unless NP = RP. We conclude the paper with a numerical analysis that guides the choice of parameters for short path algorithms and raises the possibility of super-quadratic speedups in settings that are currently beyond our theoretical analysis.
1 Introduction
1.1 Motivation
The prospect of quantum algorithmic speedups for combinatorial optimization has been heavily studied for more than two decades [FGGS00, FGG14, Has18a, Mon18, Mon20]. This interest is partially motivated by practical considerations, since combinatorial optimization problems are ubiquitous in scientific and industrial applications, and are a major source of computational bottlenecks [AAA+24, HGL+23, DMB+23]. A second principled motivation is that there are reasons to expect such a speedup, arising from the existence of quantum algorithms such as Grover’s search algorithm [Gro96], which enjoys a quadratic quantum speedup over brute force search. Since the best classical algorithms with rigorously provable runtimes for combinatorial optimization often reduce to (possibly structured) search over a large space of possible solutions, one may expect algorithms building on Grover Search to provide speedups for these algorithms. In recent years, this intuition has been largely confirmed with the development of quantum-accelerated versions that offer quadratic speedups for backtracking [Mon18, AK17] and branch-and-bound [Mon20, CMYP22], two of the main classical meta-algorithms used to obtain provable runtime guarantees. There has been some success in obtaining sub-quadratic speedups for combinatorial optimization algorithms based on Markov Chain Monte Carlo [WA08], and dynamic programming [ABI+19].
Despite this progress, there remain challenges towards leveraging quantum algorithms for combinatorial optimization. Firstly, there are many problems for which quadratic quantum speedups over the state-of-the-art classical approaches have not been demonstrated, including cases where the best algorithm is based on dynamic programming [FGS06], MCMC, and local search [Sch02, ST13]. The second challenge is more fundamental, as recent research indicates that the realization of quadratic quantum speedups is unrealistic due to constant-factor slowdowns compared to classical hardware manifesting from slower clock speeds, the overhead of error-correction, and the limited parallelizability of quantum algorithms. Realistic estimates for the resources required to execute a quantum algorithm on a scale where it can break-even with classical computing result in quantum runtimes exceeding many days. The viability of practical realization of a polynomial speedup increases with the degree of the speedup. Recent work indicating that the outlook for realizing a quartic speedup, when considering all overheads, can be much more realistic, requiring quantum runtimes on the scale of hours instead of days [BMN+21]. It is thus of fundamental importance to investigate whether it is possible to obtain “super-quadratic” speedups for combinatorial optimization.
The general quantum speedups for backtracking [Mon18], branch-and-bound [Mon20], dynamic programming [ABI+19] and MCMC methods either rely directly on Grover Search, or closely related methods like amplitude estimation, quantum minimum finding, or discrete time quantum walks. These frameworks therefore admit at most quadratic speedups by construction. Furthermore, in the case of unstructured search, backtracking and branch-and-bound, quadratic speedups can be shown to be the best one can hope for in the oracle setting. The study of super-quadratic speedups necessitates the investigation of mechanisms for quantum speedup beyond Grover Search. It is also likely that these speedups must leverage problem specific structure to circumvent the aforementioned lower bounds in the oracle setting.
A simpler, but very non-trivial question, is whether there are quantum algorithms for combinatorial optimization that achieve Super-Grover speedups. That is, a super-quadratic speedup over unstructured search, but not necessarily the best classical algorithm. An important first step towards rigorously obtaining positive results in this direction was the “short path” quantum algorithm from Hastings [Has18a], which was demonstrated to solve combinatorial optimization problems with runtime $\mathcal{O}^{\ast}\left(2^{(0.5-c(n))n}\right)$, where $c(n)$ is a positively valued function of $n$. Here and in the sequel, we use the notation $\mathcal{O}^{*}(2^{h(n)})$ to indicate an upper bound on the runtime of the form $\operatorname{poly}(n)2^{h(n)}$ that holds for sufficiently large $n$. A limitation of this result was that the rigorously established bounds on $c(n)$ asymptote to 00 as $n$ increases, leading to sub-polynomial improvements over Grover search. In a followup work that built on the framework of Hastings (but using a significantly modified algorithm and analysis), Dalzell et al. [DPCB23] gave an algorithm that obtains strictly Super-Grover speedups for several combinatorial optimization problems, i.e., the algorithm achieves a runtime $\mathcal{O}^{*}\left(2^{(0.5-c)n}\right)$, where $c$ is a positive parameter independent of $n$. Our work largely builds on the algorithm of Dalzell et al. [DPCB23], which we henceforth refer to an improved short path algorithm due to its connection with Hastings’ original work.
The successful demonstration of Super-Grover speedups leads to natural optimism that similar techniques could be used in principle to show super-quadratic speedups over the best known classical algorithm. There remain several challenges towards such a demonstration. On one hand, the speedups shown in [DPCB23] are only larger than quadratic by very small factors. On the other hand, the best performing classical algorithms for well-studied combinatorial optimization problems (while still exponential-time) are significantly faster than unstructured search. For instance, the well-known 3-SAT problem can be solved in time $\mathcal{O}^{*}\left(2^{0.39n}\right)$, the maximum independent set of an $n$-vertex graph can be found in time $\mathcal{O}^{*}(2^{0.258n})$, and the exact ground state of the Sherrington-Kirkpatrick model can be determined in time $\mathcal{O}^{*}\left(2^{0.45n}\right)$. As a consequence, the runtimes established in [DPCB23] are in most cases slower than the best classical algorithm. An exception to this is the problem of minimizing the energy of $k$-spin models, for which there has been limited study of classical algorithms. It is important to note that the mathematical analysis of [DPCB23] does not make much effort to optimize the parameters of the algorithm, and the actual runtime is predicted to be better than the theoretical predictions. However, we provide results from a numerical simulation of the algorithm in [DPCB23] which indicate that, even when the parameters are optimized, the speedup is likely to be insufficient by itself. For Maximum Independent Set on graphs with up to 21 vertices, the best runtime scaling (using a penalty term to enforce the constraints) achieved in our experiments is around $\mathcal{O}^{*}\left(2^{0.400n}\right)$. This is slower than the best classical algorithm despite optimizing the parameters, indicating that the frameworks in [Has18a, DPCB23] must be further generalized if genuine super-quadratic speedups are the goal.
A second consideration is that many combinatorial optimization problems of industrial importance are constrained, including well-studied examples such as Maximum Independent Set, Maximum/Minimum Bisection, Vertex and Set Cover, Portfolio Optimization, and Hamiltonian Cycles. The current short path algorithms can only incorporate constraints by adding penalty terms to the cost function, an approach rarely used by the state-of-the-art algorithms for combinatorial optimization. Furthermore, the runtime of the short path algorithm scales with the number of bit-strings defined on the unconstrained solution space, which can often be much larger than the number of bit-strings that satisfy all constraints in the problem formulation. As an example, consider a combinatorial optimization problem with a constraint requiring solutions to have Hamming weight $\lfloor n^{\alpha}\rfloor$ for some $10<\alpha<$ 1. The number of feasible bit-strings is $2^{(1-\alpha)n^{\alpha}\log(n)}$, which is asymptotically smaller than $2^{n}$ by a significant margin. For such constrained problems, the short path algorithm cannot even offer a super-quadratic speedup over unstructured search (if the search is restricted to the feasible region).
This paper seeks to address these limitations by analyzing a generalized version of short-path algorithms that, under some technical conditions, obtains super-quadratic speedups over classical algorithms that search the space of solutions using samples from the stationary distribution of a Markov Chain. We refer to such algorithms as Markov Chain Search, and the framework can be simply described [^1] We refer the reader to Section 2 for background on Markov Chains as follows. Suppose that our aim is to minimize a real-valued cost function $H\colon\mathcal{X}\mapsto\mathbb{R}$ for a finite set $\mathcal{X}\subset\{-1,1\}^{n}$, and let $P$ be the transition matrix of a Markov Chain that mixes to a stationary distribution $\pi$ supported on $\mathcal{X}$ in $\operatorname{poly}(n)$ steps. A Markov Chain Search algorithm using $P$ simply runs the chain to draw samples from $\pi$, and keeps track of the running minimum of the samples in terms of the cost function $H$. This minimum (and the corresponding sample) serves as an estimate of the global minimum (and minimizer) of $H$ on $\mathcal{X}$. To bound the expected runtime of this algorithm, it is sufficient to bound the expected number of samples before a global minimum is encountered. Letting $\pi^{\ast}$ denote the total probability that a sample from $\pi$ is a global minimizer of $H$, it follows that drawing $\mathcal{O}\left((\pi^{\ast})^{-1}\log(\delta^{-1})\right)$ samples from $\pi$ suffices to ensure that we encounter the global minimizer with probability at least $1-\delta$. We outline this framework in Algorithm 1.
Markov Chain Search is a natural extension of unstructured search, with several advantages. Firstly, it is often possible to classically sample from distributions over the feasible set that favor lower cost solutions. A common example is sampling from the Gibbs distribution corresponding to the cost function $H$, where a solution $z\in\mathcal{X}$ is sampled with probability proportional to $\exp(-\beta H(z))$ for some parameter $\beta$ (usually called the inverse temperature). Gibbs distributions are the stationary distributions of well known chains such as the Glauber Dynamics [Gla63] (also referred to as Metropolis sampling), and the mixing times of these Markov Chains have been extensively studied in Physics and Computer Science for decades. It is evident that in a Gibbs distribution with $\beta>0$, low cost solutions are more likely than in the uniform distribution. If the global minima has significantly lower cost than most of the ensemble, this can lead to algorithms that are polynomially faster than unstructured search. We note that Dalzell et al. [DPCB23] make reference to precisely this framework when examining the possibility of faster classical algorithms for $k$-spin problems. A second advantage of Markov Chain Search arises when the feasible set is asymptotically much smaller than $2^{n}$ as discussed previously. Clearly, if there is a chain $\mathcal{M}$ that mixes in polynomial time to the uniform distribution over $|\mathcal{X}|$, the Markov Chain Search algorithm with $\mathcal{M}$ is faster than unstructured search. Even if we can only prepare a (not necessarily uniform) distribution whose support is restricted to $\mathcal{X}$, this may result in a runtime that is substantially better than unstructured search. While the classical analysis of precise runtimes using Markov Chain Search is typically challenging, we later give explicit examples of settings where this separation from unstructured search is realized.
1.2 Contributions
Our primary contribution is the formulation of the generalized short-path algorithm and the demonstration that, under certain technical conditions, it obtains a super-quadratic speedup over unstructured search. In particular we obtain quantum runtimes of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$ where $T(n)$ is the (exponential in $n$) runtime of the classical Markov Chain Search, and $c(n)>0$ is a positive parameter. When $c(n)$ is bounded below by a constant $c>0$ we say we have a true super-quadratic speedup, and when $c(n)=o(1)$ we say we have an asymptotically decaying advantage over quadratic speedup (in the vein of Hastings’ results [Has18a, Has18b, Has19a].) We go on to discuss how these conditions may be established and demonstrate explicit results in two settings.
  1. 1.
    As an example of constrained optimization, we focus on optimization over strings of fixed Hamming Weight, using a Markov Chain based on random transpositions. In this case, Markov Chain search reduces simply to a brute force search over feasible strings, i.e., those that satisfy the Hamming weight constraint. Correspondingly, we identify conditions on the cost function for which a super-quadratic speedup using the short path framework can be obtained over brute-force search restricted to feasible states. As an explicit example, we consider the Maximum-Bisection problem. We note that the new runtime by itself is not particularly interesting as the number of feasible solutions to Maximum-Bisection (i.e., the number of $n$-bit binary strings of Hamming weight $n/2$) is smaller than $2^{n}$ by a polynomial factor and thus, the runtime for Markov Chain Search is not notably faster than unstructured search for this problem. However, applying the generalized framework allows for an algorithm that is completely restricted to feasible states, and the corresponding runtime cannot be obtained using the existing frameworks found in [Has18a, DPCB23].
  2. 2.
    Our next application is to Markov Chain search with non-uniform sampling, particularly to search with Gibbs distributions prepared using the Glauber Dynamics. This analysis yields conditions for super-quadratic speedups for problems such as the Maximum Independent Set problem, the Sherrington Kirkpatrick Model, and the Ising Model on graphs of bounded degree, for which Markov Chain search outperforms unstructured search. For the maximum independent set problem on regular graphs, we demonstrate that if the degree is a sufficiently high constant, Markov Chain search with the Glauber dynamics outperforms the best known theoretical classical algorithms for this problem. We also identify a number of cases where our analysis yields algorithms with a super-quadratic advantage over Markov Chain search where the advantage is asymptotically quadratic for large $n$ (in the vein of Hastings’ original results [Has18a]).
The following theorem informally summarizes the applications of the framework.
Theorem 1.1 (Applications of Generalized Short-Path Framework (informal)).
The following optimization problems exhibit a quantum runtime of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$, where $T(n)$ is the runtime of a classical algorithm based on Markov Chain Search, and $c>0$ quantifies the degree of improvement over quadratic search:
  1. 1.
    Maximum Bisection on random graphs with constant average degree,where the Markov Chain is the transposition walk and $c(n)=\Theta(1)$.
  2. 2.
    Maximum Independent Set and Ising Model on graphs of constant maximum degree, where the Markov Chain is the Glauber Dynamics at any temperature that permits polynomial mixing, and $c(n)=\Theta(1)$.
  3. 3.
    A generalization of Maximum Bisection where the smaller partition is constrained to have $k$ nodes (where $k\leq n/2$). The Markov Chain is the transposition walk and $c(n)=\Theta\left(1/\log(n/k)\right)$.
  4. 4.
    The Sherrington-Kirkpatrick Model where the Markov Chain is the Glauber Dynamics at any temperature that permits polynomial mixing, and $c(n)=\Theta\left(1/\log(n)\right)$.
The first two items above represent a constant improvement over quadratic speedup, whereas in the latter two, the speedup decays to 00 as $n$ increases. The Markov Chain Search considered is faster than brute force search on Maximum Independent Set for random regular graphs of sufficiently high degree, and on all instances of the other problems.
From a technical point of view, while the skeleton of the framework follows that of [DPCB23] quite closely, the intermediate conditions must all be uplifted to incorporate the Markov Chain used for the base classical algorithm, and significant care is needed to obtain the generalized results. These generalizations are crucial to leverage the two main advantages of Markov Chain search over unstructured search, the ability to sample from non-uniform distributions and distributions with support restricted to feasible solutions. As we will later show, these cannot be achieved by specializations of existing frameworks. The uplift of the conditions also clarifies some aspects of the role they play in the argument. We found that approaching the original results of [DPCB23] from this new perspective led to some additional insights, which may be implicit, but are not explicitly documented in other works. We note also that the algorithm of [DPCB23] follows directly as a special case of our generalized framework by considering the bit-flip walk (or the random walk on the vertices of the hypercube). Finally, the generalized analysis in this paper allows us to greatly simplify some of the statistical mechanics arguments made in [DPCB23], and instead rely directly on some standard results from the theory of Markov Chains and this may be of separate interest.
Our techniques also allow us to shed more light on a fundamental question about the viability of true super-quadratic speedups with the short path framework. The algorithms in this paper as well as the earlier frameworks, rely on preparing a quantum state whose overlap with the global minimizers is larger than that of some easily prepared starting state and jumping to the global minimum from that state. It is apparent that if there existed a classical algorithm to sample in polynomial time from a distribution with overlap that matches that of this intermediate state, then there is a classical algorithm that finds the global minimum only quadratically slower than the short path algorithm. The advantage over search is then essentially “dequantized” and there is no hope for true super-quadratic speedups. It is therefore important to understand to what degree classical sampling techniques can approach the overlap of the intermediate state, as discussed in [DPCB23]. Most natural sampling algorithms are based on the analysis of Markov Chains and so our framework provides a useful tool to probe this question. Our results on the Maximum Independent Set for bounded degree graphs yields some concrete evidence that the advantage of the short path framework over search cannot simply be removed by classical sampling. Specifically, we demonstrate a concrete optimization problem for which the intermediate state prepared by the short path algorithm has higher overlap than any Gibbs distribution that can be prepared in polynomial time, unless NP = RP.
1.3 Related Works
As discussed at length in the above, the primary inspiration for our work comes from the earlier short path algorithms developed in the series of works [Has18a, Has19b, Has18b, DPCB23]. We now discuss other related quantum algorithms and their connections to our results. The most closely related family of algorithms are based on discrete time quantum walks. The framework of Magniez at al [MNRS11] considers a framework that closely matches the one considered here. In [MNRS11], the authors seek to accelerate a classical algorithm, that first prepares in time $\mathrm{S}$ a sample from the stationary distribution of a Markov Chain with spectral gap $\delta$, that has $\epsilon$ overlap with some marked state that can be checked in time $\mathrm{C}$, and then runs the Markov Chain to repeatedly draw and check samples from the stationary distribution. The classical algorithm finds a marked element in time $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1}(\delta^{-1}\mathrm{U}+\mathrm{C})\right)$ where $\mathrm{U}$ is the cost of simulating one step of the chain. The quantum algorithm obtains a runtime of $\mathcal{O}\left(\mathrm{S}+\epsilon^{-1/2}(\delta^{-1/2}\mathrm{U}+\mathrm{C})\right)$.
In our setting $\epsilon$ corresponds to $\pi(E^{\ast})$ and falls exponentially, whereas $\mathrm{S,U,C},\delta^{-1}$ all grow polynomially. Applying the [MNRS11] framework, we obtain a quantum runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{0.5})$. If the conditions for our framework are met, we obtain a runtime of $\mathcal{O}^{\ast}(\pi(E^{\ast})^{0.5-c(n)})$. In this setting, our algorithm accelerates the [MNRS11] framework in a manner analogous to how [DPCB23] accelerates Grover search. Another common framework for analyzing search via quantum walk is that of Szegedy [Sze04] where the quadratic speedup is over the hitting time of a marked vertex. However, since the Hitting Time $\mathrm{HT}$ from stationarity satisfies $\epsilon^{-1}\leq\mathrm{HT}\leq\epsilon^{-1}\delta^{-1}$, we have $\mathrm{HT}=\mathcal{O}^{\ast}(\epsilon^{-1})$ in our setting where $\epsilon^{-1}$ is exponentially larger than $\delta^{-1}$, and the runtime of both quantum walk frameworks match upto polynomial factors.
Aside from the works we have already mentioned, quantum algorithms have also been successfully leveraged to obtain speedups for solving linear systems of equations [HHL09, CKS17, CGJ19], computing partition functions [HW20, CH23], estimating the volume of convex bodies [CCH+23], as well as sampling from both log-concave [CLL+22] and non-log-concave [OLMW24] distributions. We note that our work is primarily concerned with exponential time search algorithms instead of the randomized approximation schemes considered in these papers. It would be interesting to understand whether our methods can be used to obtain super-quadratic speedups for exponential time counting algorithms [GLR21]. Along the line of super-quadratic quantum speedups, there is also recent work providing positive results for Tensor Principle Component Analysis [SOKB24], and approximation algorithms for combinatorial optimization [JSW+24].
1.4 Organization
The rest of the paper is organized as follows. Section 2 introduces some basic notation, background on Markov Chains, and a refresher on the original short path algorithm. A reader familiar with Markov Chains may simply skip to Section 3, which provides a technical introduction to the generalized short path framework and our main theoretical results. This section also includes some discussion on the techniques and the consequences of these results. The following sections contain the technical proofs, with results about the general framework in Section 4 and results about the applications to specific problems in Section 5.
2 Preliminaries
We write $\log$ and $\ln$ to indicate logarithms base 2 and base $e$, respectively. We denote the $i$-th element of a vector $x\in\mathbb{C}^{n}$ by $x_{i}$, and the $ij$-th element of a matrix $A\in\mathbb{C}^{m\times n}$ by $A_{ij}$. For a vector $x\in\mathbb{C}^{n}$, the matrix $\mathrm{diag}(x)\in\mathbb{C}^{n\times n}$ takes the values of $x$ on its diagonal and zero elsewhere. We write $A\succeq 0$ ($A\succ 0$) to indicate that a matrix $A\in\mathbb{C}^{n\times n}$ is positive semidefinite (positive definite), i.e., all of its eigenvalues are nonnegative (positive). For two $m\times n$ matrices $A$ and $B$, we write $A\circ B$ to indicate their Hadamard (or, element-wise) product. Note that when we say the phrase with high probability (w.h.p. for short), we imply that a result holds asymptotically in the problem size $n$ with probability $1$.
Order estimates
We define $\mathcal{O}(\cdot)$ as
$f(x)=\mathcal{O}(g(x))\iff\exists\ell\in\mathbb{R}{},\alpha\in\mathbb{R}{}_{+},~{}\text{such that}~{}f(x)\leq\alpha g(x)\quad\forall x>\ell.$
Similarly,
$f(x)=o(g(x))\iff\exists\ell\in\mathbb{R}{},~{}\text{such that}~{}f(x)<\alpha g(x)\quad\forall x>\ell~{}\text{and}~{}\forall\alpha\in\mathbb{R}{}_{+}.$
We write $f(x)=\Omega(g(x))\iff g(x)=\mathcal{O}(f(x))$. If there exists positive constants $\alpha_{1}$ and $\alpha_{2}$ such that
$\alpha_{1}g(x)\leq f(x)\leq\alpha_{2}g(x)\quad\forall x>0,$
then we write $f(x)=\Theta(g(x))$.
We also define $\widetilde{\mathcal{O}}(f(x))=\mathcal{O}(f(x)\cdot\textup{polylog}(f(x)))$ and $\mathcal{O}^{*}(2^{f(x)})=\mathcal{O}(2^{f(x)}\cdot\operatorname{poly}(x))$.
2.1 Probability toolbox
In what follows let $\mathcal{X}$ be a finite set satisfying $|\mathcal{X}|=V$. A Markov chain $\mathcal{M}$ is a random process that defines movements between elements of $\mathcal{X}$. Transitions between states are determined according to a fixed probability distribution, and can be represented by an $V\times V$ (though not necessarily symmetric) transition matrix $P$. The entry $P_{kj}:=P(k,j)$ is the probability of making a transition from $k$ to $j$, and the rows of $P$ sum to 1 to preserve normalization $\sum_{j\in\mathcal{X}}P(k,j)=1$; we say that such a matrix is stochastic.
One step in the chain obeys
$\mu^{(t)}P=\mu^{(t+1)}\quad\forall t\geq 0.$
For any initial distribution $\mu^{(0)}\in\mathbb{R}^{n}$ over $\mathcal{X}$, the distribution after $t$ steps of the walk is
$\mu^{(0)}P^{t}=\mu^{(t)}\quad\forall t\geq 0.$
We say that a distribution $\pi$ over $\mathcal{X}$ is a stationary distribution if
$\pi P=\pi.$
For a function $f:\mathcal{X}\mapsto\mathbb{R}$, we define
[…]
and for two such functions $f,g$ their $\pi$ inner product is
$\langle f,g\rangle_{\pi}=\sum_{x\in\mathcal{X}}\pi(x)f(x)g(x).$ (2)
The reversed Markov Chain of $\mathcal{M}$ is defined as the Markov chain $\mathcal{M}^{*}=(\mathcal{X},P^{*},\pi)$, which shares the stationary distribution $\pi$ of $\mathcal{M}$, and $P^{*}$ is defined by the equation:
$\pi(x)P(x,y)=\pi(y)P^{*}(y,x).$ (3)
The chain $\mathcal{M}$ is called reversible if $P^{*}=P$.
The total variation distance between two probability distributions $\mu$ and $\nu$ on $\mathcal{X}$ is defined by
$\textup{TV}(\mu,\nu):=\frac{1}{2}\|\mu-\nu\|_{1}.$
It also satisfies the folowing variational formula
$\textup{TV}(\mu,\nu)=\textup{sup}_{f:\lVert f\rVert\leq 1}\frac{1}{2}(\mathbb{E}_{\mu}[f(x)]-\mathbb{E}_{\nu}[f(x)]).$
Another important quantity for comparing probability distributions is the Kullback–Leibler (KL) divergence (or, relative entropy) of two probability distributions $\mu$ and $\nu$ over $\mathcal{X}$:
$\operatorname{KL}(\mu\|\nu):=\sum_{x\in{\mathcal{X}}}\mu(x)\ln\left(\frac{\mu(x)}{\nu(x)}\right).$
For KL divergence we also have the Donsker and Varadhan’s variational formula [DV83] for $\textup{KL}(\mu\|\nu)$:
$\textup{KL}(\mu\|\nu)=\sup_{f\in\mathcal{F}}\left\{\mathbb{E}_{\mu}[f(x)]-\ln(\mathbb{E}_{\nu}[\exp(f(x))]\right\},$
where $\mathcal{F}$ denotes the set of all measurable functions.
The mixing time of a Markov chain is the amount of time it takes for the distance to stationarity to be small:
$t_{\text{mix}}(\varepsilon):=\min\left\{t:d(t)\leq\varepsilon\right\},$
where $d(t):=\sup_{\mu}\textup{TV}(\mu P^{t},\pi)$. The mixing time of a reversible Markov chain is related spectral properties of $P$. In this case, the matrix $P$ is similar to a symmetric matrix, and thus diagonalizable. The eigenvalues of $P$ can be ordered as
$11\geq\lambda_{1}\geq\cdots\geq\lambda_{N}\geq-$ 1 .
It is known that $\lambda_{1}=1$ and $\lambda_{2}<1$. We define the spectral gap of a reversible Markov chain to be:
$\delta:=1-\max_{\{\lambda_{2},\dots,\lambda_{N}\}}\{\lvert\lambda\rvert:\lambda\neq\pm 1\}.$
The relationship between mixing time and the spectral gap can be expressed as
$t_{\text{mix}}(\varepsilon)=\widetilde{\mathcal{O}}_{\frac{1}{\varepsilon}}\left(\delta^{-1}\right).$
A larger spectral gap therefore implies faster mixing, meaning that the Markov chain more rapidly converges to the stationary distribution.
Next we define the Discriminant matrix of a Markov chain, which is a useful tool for analyzing random walks.
Definition 2.1 (Discriminant matrix).
For a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$, the discriminant matrix is the operator with elements
$D(P)_{ji}:=\sqrt{P_{ij}\circ P^{*}_{ji}}=\left(\mathrm{diag}(\pi)^{1/2}P\mathrm{diag}(\pi)^{-1/2}\right)_{ij}.$ (4)
Furthermore, if $P$ is reversible, then $D(P)$ is symmetric, and the following hold:
  1. 1.
    The unique, maximum eigenvalue eigenstate of $D(P)$ is $|\sqrt{\pi}\rangle$ with eigenvalue $1$.
  2. 2.
    The spectral gap of $-D(P)$ (and equivalently $P$) is
    $\delta:=1-\max\{\lvert\lambda\rvert:\lambda\in\sigma(D(P)),\lambda\neq\pm 1\}.$ (5)
  3. 3.
    $\lVert D(P)\rVert_{2}=1$.
  4. 4.
    If $P$ is symmetric, then $D(P)=P$.
Definition 2.2 (Markov functionals).
Let $f:\mathcal{X}\mapsto\mathbb{R}$. The Dirichlet form, $\mathcal{D}(f,f)$, generated by a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is defined by
$\mathcal{D}(f,f):=\langle f,(I-P)f\rangle_{\pi},$
and if $P$ is reverisble, then $I-P$ is symmetric with respect to the $\pi$-inner product, and so $\mathcal{D}$ extends to an inner product for functions $f,g$:
[…]
The $\pi$-Variance of $f$ is defined as
$\textup{Var}_{\pi}(f):=\mathbb{E}_{\pi}[f^{2}]-(\mathbb{E}_{\pi}[f])^{2},$
and the $\pi$-Entropy of $f$ is defined as
$\textup{Ent}_{\pi}(f):=\mathbb{E}_{\pi}[f\ln(f)]-\mathbb{E}_{\pi}[f\ln(\mathbb{E}[f])].$
One can relate the variance to the Dirichlet form and the spectral gap using a Poincaré inequality:
Definition 2.3 (Poincaré inequality).
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a Poincaré inequality with constant $\delta$ if
$\mathcal{D}(f,f)\geq\delta\textup{Var}_{\pi}(f).$
For reversible Markov chains, the Poincaré constant is equal to the spectral gap. It is possible to obtain better bounds on the mixing time of a Markov chain using the so-called logarithmic Sobolev inequalities.
Definition 2.4 (log-Sobolev inequality).
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a log-Sobolev inequality with constant $\omega:=\omega_{LS}$ if
$\mathcal{D}(f,f)\geq\omega\textup{Ent}_{\pi}(f^{2}).$
Definition 2.5 (modified log-Sobolev inequality).
A Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ satisfies a log-Sobolev inequality with constant $\omega_{\textup{MLS}}$ if
$\mathcal{D}(f,\ln f)\geq\omega_{\text{MLS}}\textup{Ent}_{\pi}(f).$
The following chain of inequalities is well-known:
$\delta\geq\omega_{\textup{MLS}}\geq\omega_{\textup{LS}}.$ (6)
The Poincaré and log-Sobolev inequalities belong to a group known as the functional inequalities. They are well-defined for non-reversible chains, although Poincaré now bounds the singular-value gap, and enable one to bound the corresponding mixing time [Cha23].
Finally, we define the $P$-pseudo Lipschitz norm, which measures the smoothness of a function with respect to the transition probabilities of a Markov chain.
Definition 2.6 ($P$-pseudo Lipschitz norm).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The $P$-pseudo Lipschitz norm of $f:\mathcal{X}\mapsto\mathbb{R}$ is defined to be
[…]
2.2 Short path algorithms
Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a classical cost function satisfying $\sum_{z}H(z)=0$ for $H$ with no constant term. Consider the combinatorial optimization problem
$E^{\ast}:=\min_{z\in\left\{-1,1\right\}^{n}}H(z),$
where $E^{\ast}$ is the optimal objective value. Let $\Pi^{\ast}$ denote the orthogonal projector onto subspace spanned by optimal assignments $|z^{\ast}\rangle$.
Let $X=\sum_{i\in[n]}X_{i}$ be the transverse-field operator, where $X_{i}$ denotes the Pauli-$X$ operator acting on qubit $i\in[n]$. A well-known approach to determine some $|z^{\ast}\rangle$ is the quantum adiabatic algorithm (QAA) [FGGS00]. The QAA finds a $|z^{\ast}\rangle$ by considering the adiabatic time evolution of
$H_{b}^{(\textup{QAA})}=-(1-b)X+bH,$
as $b$ is tuned from $b=0$ to $b=1$. However, QAA is known to suffer from certain “localization” issues, which can be viewed as a quantum analogue of getting trapped in local minima, and can result in run times that are exponentially worse than classical brute-force search [AKR10]. This manifests as a result of the “avoided crossing” phenomenon, or first-order phase transition that can lead to exponentially (or even super-exponentially) small spectral gaps.
Recently, Hastings [Has18a] and Dalzell et al. [DPCB23] proposed a new paradigm for avoiding the first-order phase transition problem with the QAA. Following the approach of [DPCB23], prototypical adiabatic optimization is eschewed through two modifications. First, the term $H$ is replaced with $g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$ for a piecewise-linear function $g_{\eta}:[-1,\infty)\mapsto[-1,0]$ parameterized by $\eta\in[0,1)$:
$g_{\eta}(x):=\min\left(0,\frac{x+1-\eta}{\eta}\right),$
leading to the Hamiltonian:
$H_{b}=-\frac{X}{n}+bg_{\eta}\left(\frac{H}{|E^{\ast}|}\right),$ (7)
where $X$ has been normalized by its spectral norm, and $b\in[0,\infty)$. Second, rather than slowly evolve from $-\frac{X}{n}$ to $\frac{H}{|E^{\ast}|}$ as in the QAA, we jump from $-\frac{X}{n}$ to the ground state |$\ket{\psi_{b}}$ of $H_{b}$ for some value of $b$ that is independent of $n$ (where the spectral gap is guaranteed to be large), and then jump from $H_{b}$ to the ground-state space of $\frac{H}{|E^{\ast}|}$. Note in [DPCB23] they also allow for scaling $H$ by an overestimate of $\lvert E^{*}\rvert$, for simplicity we just stick with scaling by $\lvert E^{*}\rvert$.
The jumps are accomplished using a unitary $U$, which combines phase estimation and amplitude amplification. For a high-level understanding, suppose we seek to enact a jump between two Hamiltonians $H_{1}$ and $H_{2}$, each acting on $n$ qubits. Denote the ground state of $H_{1}$ by |$\ket{\psi_{1}}$, and let $\Pi_{2}$ be the projector onto the ground space of $H_{2}$. The unitary $U$ first employs phase estimation to implement reflection operators $R_{1}$ and $R_{2}$ that reflect about the state |$\ket{\psi_{1}}$, and the groundspace of $H_{2}$, respectively. If $\delta_{j}$ is the spectral gap of Hamiltonian $H_{j}$, the operator $R_{j}$ can be implemented up to error $\varepsilon$ using $\delta_{j}^{-1}\log(1/\varepsilon)$ calls to a block-encoding of $H_{j}$, and often realizable using $\operatorname{poly}(n)$ gates. When $H_{j}$ is a classical Hamiltonian, $R_{j}$ can be implemented using $\operatorname{poly}(n)$ gates irrespective of $\delta_{j}$. From here, the unitary $U$ employs fixed-point amplitude amplification [YLC14] to implement $\frac{\Pi_{2}\ket{\psi_{1}}}{\left\|\Pi_{2}\ket{\psi_{1}}\right\|}$, requiring at most $\mathcal{O}\left(\left\|\Pi_{2}\ket{\psi_{1}}\right\|^{-1}\log(1/\varepsilon)\right)$ applications each of $R_{1}$ and $R_{2}$.
The algorithm is initialized to |+:=|$\ket{\boldsymbol{+}}:=\ket{+}^{\otimes n}$, the ground state of $-\frac{X}{n}$. Then, the ground state |$\ket{\psi_{b}}$ of $H_{b}$ is prepared by jumping from $-X/n$ to $H_{b}$. Finally, we prepare $\frac{\Pi^{\ast}\ket{\psi_{b}}}{\lVert\Pi^{\ast}\ket{\psi_{b}}\rVert}$ by jumping from $H_{b}$ to the classical Hamiltonian $\frac{H}{{\left|E^{\ast}\right|}}$. The state $\Pi^{\ast}\ket{\psi_{b}}$ is a superposition of optimal solutions to $\min_{z\in\{\pm 1\}^{n}}H(z)$, and thus measurement in the computational basis yields an optimal bit-string $z^{\ast}$ with high probability. The first jump is small (in the sense that the success probability is nearly 1), whereas the second jump is large (the success probability is exponentially small). The resulting time complexity scales as $\mathcal{O}^{*}\left(2^{\left(\frac{1}{2}-c\right)n}\right)$, indicating a super-Grover speedup when $c>0$. In [Has18a], the order of short and long jump steps is reversed. We refer to both approaches [Has18a, DPCB23] as short path algorithms.
3 Technical Overview
3.1 Framework
Define $\mathcal{X}\subseteq\{-1,1\}^{n}$ and let $H\colon\mathcal{X}\mapsto\mathbb{R}$ be a cost function. We are interested in (exactly) determining $z^{\ast}\in\mathcal{X}$ such that
$z^{\ast}\in\operatorname*{arg\,min}_{z\in\mathcal{X}}H(z).$
We also use $H$ to refer to a diagonal Hamiltonian in a Hilbert space with a basis indexed by $z\in\mathcal{X}$, with $\langle z|H|z\rangle$ for the Hamiltonian identified with $H(z)$. Whether we are referring to the quantum Hamiltonian or the function will be clear from context. We define $E^{\ast}:=\min_{z\in\mathcal{X}}H(z)$ and assume everywhere that the cost is scaled to ensure $E^{\ast}<0$. We will further assume that $|\mathcal{X}|$ is super-polynomial, as our primary concern is with super-quadratic speedups over exponential time algorithms. If $\pi$ is a distribution such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer and there exists a Markov Chain with transition matrix $P$ that mixes to stationary distribution $\pi$ such the mixing time is bounded by $t_{\mathrm{mix}}(\varepsilon)=\operatorname{poly}(n,\log(\varepsilon^{-1}))$.
Algorithm 1 MarkovChainSearch
Prerequisites: Solution space $\mathcal{X}\subset\{-1,1\}^{n}$, Cost function $H\colon\mathcal{X}\mapsto\mathbb{R}$, distribution $\pi$ such that $\pi(E^{\ast})$ is the probability that a sample from $\pi$ is a global minimizer, a Markov Chain with transition matrix $P$ and mixing time $t_{\mathrm{mix}}(\pi(E^{\ast})/2)=\mathcal{O}(\operatorname{poly}(n))$.
      Input: Description of $\operatorname{poly}(n)$ time procedures to evaluate $H$ and perform a step of the Markov Chain described by $P$, failure probability $\delta$.
      Output: A global minimizer $z^{\ast}$ of $H(z)$ over $\mathcal{X}$.
1:Set $i=0$, $z^{(0)}$ to an arbitrary point in $\mathcal{X}$.
2:while $i\leq\frac{2}{\pi(E^{\ast})}\log\left(\frac{1}{\delta}\right)$ do
3:     Simulate $t_{\mathrm{mix}}(n)$ steps of $P$ to obtain sample $\tilde{z}$
4:     if $H(\tilde{z})\leq H(z^{(i)})$ then
5:         Set $z^{(i+1)}=\tilde{z}$ and $i\leftarrow i+1$.
6:     else
7:         Set $z^{(i+1)}=z^{(i)}$ and $i\leftarrow i+1$.
8:Output $z^{(i)}$.
Under these conditions, it follows that Algorithm 1 finds the global minimizer of $H$ in time $\mathcal{O}^{\ast}\left((\pi(E^{\ast}))^{-1}\right)$. Our framework seeks to accelerate this runtime, and so we assume the same setting for the quantum algorithm. In order to define the quantum framework we must access $H,P,$ and $\pi$. We assume the existence of the following subroutines:
Assumption 1 (Quantum Input Assumptions).
The following subroutines are used as primitives in the Generalized Quantum Short Path framework.
  1. 1.
    Initial State Preparation: We assume the existence of a unitary $U_{\pi}$ implementable using $\operatorname{poly}(n)$ gates, such that $U_{\pi}|0\rangle=|\sqrt{\pi}\rangle\coloneqq\sum_{z\in\mathcal{X}}\sqrt{\pi(z)}|z\rangle$.
  2. 2.
    Block-encoding of Markov Chain: Suppose that $P$ is the transition matrix of a reversible Markov chain, then the discriminant operator $D(P)$ (see Definition 2.1) is Hermitian. We assume the existence of a unitary $U_{D(P)}$, implementable with $\textup{poly}(n)$ gates, that is a $(\kappa_{1},a)$ block-encoding of $D(P)$ for $\kappa_{1}=\mathcal{O}(\operatorname{poly}(n))$.
  3. 3.
    Block-encoding of Cost Function: We assume the existence of a unitary $U_{H}$, implementable with $\textup{poly}(n)$ gates, that is a $(\kappa_{2},a)$ block-encoding of $H$ for $\kappa_{2}~{}=~{}\mathcal{O}(\operatorname{poly}(n))$.
We will justify Assumption 1 for each application of the framework. Note that preparing a block-encoding of the cost function is straightforward given a $\operatorname{poly}(n)$ size circuit to evaluate it at a single point, and we do not make this analysis in every case. We also do not explicitly mention these input assumptions in each of our results to avoid cluttering the presentation, but they are prerequisites for the input model in each case. With this setup, we can define the generalized short path framework. We define a generalized “short path Hamiltonian” $H_{b}$ as
$H_{b}=-D(P)+bg_{\eta}\left(\frac{H}{|E^{\ast}|}\right),$ (8)
where $D(P)$ is the Discriminant matrix corresponding to $P$ and $g_{\eta}$ is defined similarly to [DPCB23], by
$g_{\eta}(x):=\min\left(0,\frac{x+1-\eta}{\eta}\right).$
The block-encoding for $H_{b}$ can be constructed using the linear-combination-of-block-encodings technique [GSLW19] using $U_{D(P)}$ and $U_{H}$.
The framework is specified in Algorithm 2. The implementation of the jumps uses the framework from [DPCB23, Proposition 21] that performs fixed point amplitude amplification using reflections constructed from the block-encodings of the Hamiltonians and the Quantum Singular Value Transform. The overall runtime of Algorithm 2, in terms of the number of queries to $U_{\pi}$ and block-encodings of $H$ and $D(P)$, is
$[\min(\text{Gap}(-D(P)),\text{Gap}(H_{b}))]^{-1}\left(\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert^{-1}+\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}\right),$ (9)
where $\Pi^{\ast}$ is the projector onto the ground subspace of $H$. As in previous papers, we refer to one of the steps in the algorithm as the short jump and another as the long jump. The reason for this is that the choices of $b,\eta$ made for applications of the framework will always ensure that the short jump can be carried out with a polynomial number of queries to $U_{\pi},U_{D(P)},$ and a block-encoding of $H_{b}$. Thus when including Assumption 1, the runtime of the algorithm is therefore primarily determined by the long jump, and under the appropriate conditions is bounded by $\mathcal{O}^{\ast}(\pi(E^{\ast})^{-(0.5-c)})$, leading to a super-quadratic speedup over Markov Chain Search.
Algorithm 2 GeneralizedShortPathAlgorithm
Input: Algorithmic parameters $b,\eta$, Problem parameters $H,P,\pi,E^{\ast}$, which define $H_{b}$ in Equation (8).
      Output: an optimal assignment $z^{\ast}$ for $H$.
1:Prepare |$\ket{\sqrt{\pi}}$, the ground state of $-D(P)$.
2:Short Jump: Prepare |$\ket{\psi_{b}}$ up to exponentially small error with jump $-D(P)\rightarrow H_{b}$.
3:Long Jump: Prepare $\frac{\Pi^{\ast}\ket{\psi_{b}}}{\lVert\Pi^{\ast}\ket{\psi_{b}}\rVert}$ up to exponentially small error with jump $H_{b}\rightarrow\frac{H}{{\left|E^{\ast}\right|}}$.
Our bounds on the runtime rely on two conditions, that we view as uplifted versions of corresponding notions in [DPCB23] to the case of general Markov Chains. Our first conditons captures the smoothness of the cost function under applications of the transition matrix.
Definition 3.1 ($\Delta_{P}$ stability).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. We say that the cost Hamltonian $H$ is $\Delta_{P}(\eta)$ stable under $\mathcal{M}$ if
[…]
where $h_{\eta}(x):=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$. If the conditions holds for all $10<\eta<$ 1 we omit it and simply say $H$ is $\Delta_{P}$ stable under $\mathcal{M}$.
The analysis in the following sections will clarify that $\Delta_{P}$-stable is a generalization of the $\alpha$-subdepolarizing condition introduced in [DPCB23]. In fact, it is equivalent to a more syntactically obvious generalization of the $\alpha$-subdepolarizing condition, we state Definition 3.1 as the primary condition as it is easier to demonstrate and interpret in most cases.
The next conditon is a generalization of the spectral density condition of [DPCB23]. We capture the idea that the measure (according to $\pi$) of the set of solutions $z$ for which $g_{\eta}\left(\frac{H(z)}{|E^{\ast}|}\right)\neq 0$ is polynomially related to the measure of the global minimizer, for some value of $\eta$. In other words, sampling from $\pi$ does not allow one to approximately minimize $H$ to arbitrary constant relative error, super-polynomially faster than finding the exact minimum. If this condition is violated, the problem admits a simple classical sub-exponential time approximation scheme. We define this condition as follows:
Definition 3.2 ($\gamma$ Spectral Density).
The cost Hamiltonian $H$ is said to satisfy the $\gamma$ spectral density condition with respect to the stationary distribution $\pi$ if:
$\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma}.$
We are now ready to state the main results concerning our framework. These results are subject to further technical conditions on the Markov Chain used for the search, and are formulated in terms of conditions that lead to efficient mixing time bounds. We have two variants of our result, the first relies on a logarithmic-Sobolev inequality for $P$. We have the following result:
Theorem 3.3 (informal).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, that satisfies the $\gamma$ spectral density condition for some parameters $\eta,\gamma$. In addition, suppose $\mathcal{M}$ satisfies an $\omega$ log-Sobolev inequality. If $\omega^{-1}=\Theta(\ln(1/\pi(E^{\ast}))$, then there exists a constant $b$, such that under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\omega^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right).$
We also present a variant of the above result that only relies on the weaker Poincaré inequality.
Theorem 3.4 (informal).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversbile Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, that satisfies a spectral density condition. In addition, suppose $\mathcal{M}$ satisfies a $\delta$ Poincaré inequality. If $\delta$ is independent of the problem size, then there exists a constant $b$, such that under Assumption 1, Algorithm 2 determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\delta^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right).$
The spectral density condition with respect to non-uniform starting states presents a technical challenge, since unlike the uniform distribution over all bitstrings of length $n$, they are no longer product measures. Fortunately, the condition that $\pi$ is the ground state of a fast mixing Markov Chain allows for simplification via concentration inequalities for Markov chains, e.g. the so-called Herbst argument [Lal13], that allows the spectral density conditions to be established as long as the cost function satisfies has an appropriately bounded pseudo-Lipschitz norm $\lVert H\rVert_{P}$. Theorems 3.3 and 3.4 are proven in Section 4, in order to establish specializations that rely on a bounded pseudo-Lipschitz norm.
3.2 Applications
The fact that our main results in Theorems 3.3 and 3.4 rely on functional inequalities implies many possibilities for interesting algorithmic speedups. Once a Markov Chain with the right properties (log-Sobolev or Poincaré with the proper parameters) is identified, we can derive conditions on cost functions that for which we have super-quadratic speedup over Markov Chains. As a simple toy example, consider an expander graph of size $2^{n}$, where we are given access to a polynomial time oracle that outputs the edges incident on any vertex. Since the graph is an expander, the graph random walk satisfies a Poincaré inequality with constant $\delta$. Consider any assignment of costs to the nodes such that the difference in cost between any two endpoints of an edge is bounded above by a constant. It follows from our results, that the generalized short path framework can find the node with minimum cost with super-quadratically fewer queries than searching with the random walk on the graph. A systematic study of cost functions that yield a speedup for various Markov Chains may lead to some interesting insights. For this paper, however, we focus on identifying connections to problems of general interest.
3.2.1 Optimization with Fixed Hamming Weight
We first consider optimization problems for which feasible solutions are bitstrings of fixed Hamming Weight. A well-studied example of this setting is Max-Bisection [FJ97, DMS17], defined as follows
$\mathcal{C}^{*}_{\frac{n}{2}}:=\min_{x\in\{-1,1\}^{n}}\left\{-\frac{1}{2}\sum_{i<j}e_{ij}(1-x_{i}x_{j}):\lvert x\rvert=\frac{n}{2}\right\},$ (MaxBisection)
The algorithm of [DPCB23] does not directly yield useful results for such a problem. Firstly, the framework does not naturally incorporate constraints. More importantly, although one could attempt to enforce the constraints by means of penalty terms, this prohibits the possibility of super-Grover speedups. To see why this is the case, recall that the algorithm of [DPCB23] is simply our Algorithm 2 with the Markov Chain chosen to be the random walk on the edges of the hypercube. Since every transition of such a walk changes the hypercube, the best possible value of $\Delta_{P}$ for the stability condition to be satisfied is of the same order as the penalty terms. On the other hand, the penalty terms must be of the same order as the cost function in order to guarantee that constraints are satisfied. By inspection of Theorem 3.3 we observe that no super-Grover speedup is possible via the penalty-based approach (more details on the penalty approach are in Appendix C). We overcome this challenge by employing a generalized framework that uses a Markov Chain, specifically the Bernoulli-Laplace diffusion or transposition walk, which preserves Hamming weights and transitions from a starting string of weight $k$ to the equal superposition over all such strings. In Section 5.1, we present a condition on cost functions over Hamming Weight Slices $k$ for which we obtain runtimes of the form $\operatorname{poly}(n)\binom{n}{k}^{0.5-c}$ for a constant $c$.
We study a generalization of MaxBisection which we term MaxCut-Hamming, defined as
$\mathcal{C}^{*}_{k}:=\min_{x\in\{-1,1\}^{n}}\left\{-\frac{1}{2}\sum_{i<j}e_{ij}(1-x_{i}x_{j}):\lvert x\rvert=k\right\}.$ (MaxCut-Hamming)
In Section 5.1 we prove that the Generalized Short Path Framework achieves an overall runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c}\right)$ for MaxCut-Hamming on Erdős-Rényi random graphs when $k=\Theta(n)$ (which includes MaxBisection as a special case). We also demonstrate that under the assumption of spectral density, our approach achieves a runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c(n)}\right)$ with $c(n)=\Theta\left(\log(n/k)^{-1}\right)$, and thus the super-quadratic advantage over Markov Chain search decays as $n\to\infty$ (similar to the results of  [Has18a]).
3.2.2 Glauber Dynamics
Glauber Dynamics [Gla63] is a well known sampling algorithm designed to sample from the Gibbs measures corresponding to Hamiltonians such as the Ising or Hardcore models. Since sampling from a Gibbs measure at arbitrarily high inverse temperatures is equivalent to exact optimization, for most hard problems there exists a critical threshold beyond which the Glauber dynamics no longer mixes efficiently. Performing Markov Chain search with Glauber dynamics has two advantages, if the problem is constrained then it provides a natural way to search with a distribution whose support is restricted to feasible solutions only. On the other hand, if the Glauber dynamics mixes for positive inverse temperatures, then low-energy solutions are disproportionately favored compared to the uniform distribution and the result in Markov Chain Search is asymptotically faster than unstructured search (see Lemma 5.20. In each case, we will consider classical algorithms that search using the Glauber dynamics at an inverse temperature slightly below the critical threshold where mixing takes exponential time. We demonstrate in Section 5.2 that for three models of interest, we obtain super-quadratic speedups over polynomially-mixing Glauber dynamics. These models are:
  1. 1.
    The Maximum Independent Set problem (or hardcore model) on graphs of constant maximum degree: For this problem the Glauber dynamics is shown to mix only upto critical temperatures that are negative. This means that our starting distribution favors small sets compared to large sets. However, there is the advantage that the Gibbs distribution has support only on independent sets (which are usually much fewer in number than $2^{n}$, which is the total number of subsets). In the case of random regular graphs of sufficiently high degree, we show that this Markov Chain Search algorithm is faster than unstructured search as well as the best known combinatorial algorithms for Maximum Independent Set.
  2. 2.
    The Ising Model on random regular graphs of constant maximum degree: The Ising Model is an unconstrained optimization problem and there are $2^{n}$ feasible solutions. In this setting, the Glauber dynamics mixes upto a positive critical inverse-temperature, thus the starting stationary distribution favors low energy solutions and the Markov Chain search is faster than unstructured search.
  3. 3.
    The Sherrington Kirkpatrick Model: Like the Ising Model, the Sherringtok Kirkpatrick model is also unconstrained and the Glauber dynamics mixes upto a positive inverse temperature. In this case, however, the exponent of our advantage over quadratic speedup follows with $n$. Specifically, we show a quantum runtime of $\mathcal{O}^{\ast}\left((\pi(E^{\ast})^{0.5-c(n)}\right)$ where $c(n)=\Theta(1/\log(n))$.
3.2.3 Super-Quadratic Speedup Over any Polynomial Time Gibbs Sampler
Our result regarding the Maximum Independent Set (or, hardcore model) on graphs of bounded degree allows us to go a step further and argue that it is very likely that the generalized short path can achieve a super-Grover speedup over all polynomial time Gibbs samplers (whether or not they are based on Gibbs sampling). The key observation is that for the hardcore model there is a critical fugacity $\lambda_{c}(d)=\exp(-\beta_{c}(d))=\frac{(d-1)^{d-1}}{(d-2)^{d-2}}$ such that for graphs of maximum degree $d$, Glauber dynamics mixes in time $\mathcal{O}(n\log(n))$ for any $\lambda<\lambda_{c}(d)$. However, it has been shown that computing the partition function of the hardcore model is NP-hard for any fugacity $\lambda>\lambda_{c}(d)$ [Sly10, SS14]. Due to the well known reduction between sampling and counting, it must therefore be NP-hard to sample from the corresponding Gibbs distribution. We therefore have an almost complete classification of the Gibbs measures corresponding to the hardcore model at different fugacity: it is either NP-hard to sample from the distribution, or the Glauber dynamics mixes in polynomial time. Our results in Section 5.2.1 establish that in the latter setting, the “short jump” of Algorithm 2 takes only polynomial time and produces a state $|\psi_{b}\rangle$ such that it is NP-hard to sample from any Gibbs distribution $\pi_{\beta}$ obeying $\pi_{\beta}(z^{\ast})=\Omega\left(|\langle\psi_{b}|z^{\ast}\rangle|^{2}\right)$.
Consequently, there exists a quantitative separation between the ground states of the short path Hamiltonian and Gibbs distributions. It is notable that the Markov Chain methods are obstructed at the critical fugacity due to the development of long range correlations. Such long range correlations in the ground state are also likely to lead to a vanishing gap for the short path Hamiltonian. This indicates that the ground state of the short path algorithm increases the overlap with the ground state, without creating long range correlations. Further understanding of the qualitative differences between the ground state and classical measures such as the Gibbs distribution, may shed light on the mechanisms of the short path algorithm.
We note that we resort to computational assumptions above only to show that the ground state $|\psi_{b}\rangle$ has higher overlap with the optimal solution than an efficiently sampleable Gibbs distribution. It follows unconditionally from our analysis that the ground state itself does not encode a Gibbs distribution beyond the critical mixing threshold (or indeed any distribution using which we can sample from such a Gibbs distribution by rejection sampling). In particular, Lemma 4.20 shows that for any ground state prepared by a short jump that is covered by our analysis, the trace distance from the starting state, and hence the total variation distance from the starting distribution is exponentially small.
4 Generalized Short Path Framework
This section provides a simplified and generalized analysis of the short path algorithm presented in [DPCB23]. It also highlights limitations of the current method of analysis, and describes a general recipe for the determining a speedup over Grover.
4.1 Summary of main results
In what follows, let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain over a finite set $\mathcal{X}$. Here $\pi$ denotes a stationary distribution that is uniform over $\mathcal{X}$ and $P$. It is assumed that the spectral gap of $P$ is lower-bounded by $\delta$. Using the discriminant matrix of $P$, we can define a more general short-path Hamiltonian $H_{b}$ that allows one to work with mixing operators other than $-\frac{X}{n}$.
Definition 4.1 (Short-path Hamiltonian $H_{b}$).
Consider a reversible Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. Let $H:\mathcal{X}\mapsto\mathbb{R}$ be a cost Hamiltonian. The short-path Hamiltonian $H_{b}$ is given by
$H_{b}:=-D(P)+bg_{\eta}\left(\frac{H}{|E^{\ast}|}\right),$
where $D(P)$ is the discriminant matrix of $P$, and
$g_{\eta}(x):=\min\left(0,\frac{x+1-\eta}{\eta}\right).$
More generally $g_{\eta}:[-1,\infty)\mapsto[-1,0]$ can be a non-decreasing, concave function that is differentiable at every point where it is non-zero. However, the specific choice we make is sufficient for our purposes. We will also sometimes refer use the notation $G_{\eta}:=g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$.
One major component of the analysis of Algorithm 2 is determining an upper bound on $b$ for which the spectral gap of the short-path Hamiltonian $H_{b}$ is still large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. This upper bound serves as a proxy for how large the “short jump” is. A second major component is determining the increased overlap with the optimal solution provided by the short-jump. To do so, we rely on the definition of $\Delta_{P}$ stability, which we restate below.
Definition 4.2 ($\Delta_{P}$ stability).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. We say that the cost Hamiltonian $H$ is $\Delta_{P}(\eta)$ stable under $\mathcal{M}$ if
[…]
where $h_{\eta}(x):=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$.
If the short jump can be accomplished efficiently, then $\Delta_{P}$ stability captures whether the short path approach provides a super-Grover runtime. This condition also has an intuitive interpretation. If we consider the optimization landscape defined by $\mathcal{M}$, $H$ and a well (controlled by $\eta$) around the global minimum with energy $E^{*}$, then we do not want the energy to increase too much when moving within and around the well. Specifically, for a super-Grover runtime it should hold that $\Theta\left(\frac{\lvert E^{*}\rvert}{\ln(1/\pi(E^{*}))}\right)$, where $\pi$ is the stationary distribution of $\mathcal{M}$. It is worth remarking that if $\eta=0$, we recover the quantum unstructured search algorithm.
It turns out that any upper bound on $\Delta_{P}(\eta)$ suffices when bounding the runtime. For example, it is a simple consequence of Jensen’s inequality that we can take $\Delta_{P}(\eta)$ to be $\sqrt{\lVert\psi\rVert_{P}}$ with $\psi=H$. A key technical contribution of this work is to reduce the conditions for determining whether a super-Grover runtime is possible to determining the log-Sobolev constant $\omega$ and the $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$ (or $\Delta_{P}$) for cost Hamiltonian $H$.
We summarize our main result below:
Theorem 4.3.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a log-Sobolev constant $\omega$. If $b$ satisfies
$b<b^{\ast}:=\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right),$
where
$\gamma=\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))},$
then there exists a short-path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\omega^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right).$
Note that any upper bound on $\Delta_{P}$ suffices, for example one may use $\sqrt{\lVert H\rVert_{P}}$.
Proof.
The proof is evident after combining the statements of Theorems 4.28, 4.15, Lemma 4.25, and Corollary 4.9
We also present a variant of the above result that only relies on a Poincaré inequality.
Theorem 4.4.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a Poincaré constant $\delta$. If $b$ satisfies
$b<b^{\ast}:=\delta\frac{4\sqrt{6}-1}{10},$
then there exists a short-path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
$\mathcal{O}\left(\textup{poly}(n)\delta^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln(1/\pi(E^{\ast}))\Delta_{P}}\right)}\right).$
Note that any upper bound on $\Delta_{P}$ suffices, for example one may use $\sqrt{\lVert H\rVert_{P}}$.
Proof.
The proof is evident after combining the statements of Theorems 4.28, 4.17, Lemma 4.25, and Corollary 4.10. ∎
In general, the log-Sobolev constant $\omega$ can be significantly smaller than the spectral gap of the chain $\delta$, however, we argue that this is not the case when Theorem 4.3 provides a super-Grover runtime. Specifically, Theorem 4.3 requires that $b^{*}$ is a constant, and by extension, implies we need $\omega^{-1}=\Theta(\ln(1/\pi(E^{*})))$. For example, for a very hard problem, where Markov Chain search finds an optimal assignment with exponentially-small probability, i.e., $\Theta(\ln(1/\pi(E^{*})))=\Theta(n)$, the condition on $b^{*}$ will imply that $\omega$ will be large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. Thus, in cases where it provides a super-Grover runtime, Theorem 4.3 asserts that we do not get a slower runtime by using $\omega$ instead of $\delta$.
We have the following evident corollary of the above results.
Corollary 4.5.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}$. Suppose the pair $(\mathcal{M},H)$ result in $b^{\ast}$ and $\gamma$ that are independent of $n$. If
$\frac{\lvert E^{\ast}\rvert}{\Delta_{P}}=\Theta\left(\ln\left(1/\pi(E^{\ast})\right)\right),$
then under Assumption 1 there exists a short-path algorithm that finds the minimizer of $H$ over $\mathcal{X}$ with super-Grover running time. The advantage over Grover is:
$\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b^{\ast}}{\Delta_{P}}.$
4.2 Constructing Short Path Algorithms from Markov Chains
This subsection details how the results from [DPCB23] can be generalized to the setting of reversible Markov Chains. One of the main conditions from the aformentioned paper is that there should be a small number of low-energy states, effectively capturing that the underlying problem is hard. This is made precise through the spectral density condition, which we restate below.
Definition 4.6 ($\gamma$ Spectral Density).
The cost Hamiltonian $H$ is said to satisfy the $\gamma$ spectral density condition with respect to the stationary distribution $\pi$ if:
$\pi(E\leq(1-\eta)E^{\ast})\leq\pi(E^{\ast})^{\gamma}.$
The tail bound given by the spectral density condition is implied by Pseudo Lipschitzness together with a functional inequality:
Theorem 4.7 (Herbst’s Argument, adapted from Theorem 4.3 in [Lal13] ).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with log-Sobolev constant $\omega$, and $f:\mathcal{X}\mapsto\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
$\mathbb{P}_{\pi}[f\geq\mathbb{E}_{\pi}[f]+t]\leq e^{-\frac{\omega}{\lVert f\rVert_{P}}t^{2}}.$
Theorem 4.8 (Adapted from Theorem 3.5 in [Lal13]).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with Poincaré constant $\delta$, and $f:\mathcal{X}\mapsto\mathbb{R}$ is $\lVert f\rVert_{P}$ pseudo-Lipschitz. Then,
$\mathbb{P}_{\pi}[f\geq\mathbb{E}_{\pi}[f]+t]\leq e^{-\frac{\sqrt{\delta}}{\sqrt{\lVert f\rVert_{P}}}t}.$
The following corollary is immediate, and reduces spectral density to pseudo Lipschitzness and a functional inequality.
Corollary 4.9.
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with log-Sobolev constant $\omega$, and that the cost function $H$ is $\lVert H\rVert_{P}$ pseudo Lipschitz. Then,
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))}},$
with
$\gamma=\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))}.$
Corollary 4.10.
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain with Poincaré constant $\delta$, and that the cost function $H$ is $\lVert H\rVert_{P}$ pseudo Lipschitz. Then,
$\mathbb{P}_{\pi}[H\leq(1-\eta)E^{\ast}]\leq\pi(E^{\ast})^{\frac{\sqrt{\delta}((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))}},$
with
$\gamma=\frac{\sqrt{\delta}((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])}{\sqrt{\lVert H\rVert_{P}}\ln(1/\pi(E^{\ast}))}.$
As noted in previous papers, the spectral density condition is relatively weak for hard problems. For example, suppose that for any constants $\gamma$ and $\eta$ it were not satisfied, but $\pi(E^{*})=\mathcal{O}(2^{-cn})$. Then Markov Chain search can prepare an $\eta$ relative error, for $\eta$ arbitrarly close to $1$, approximate minimizer in time subexponential in $n$.
4.2.1 The Short Jump
The short jump is defined as the preparation of $|\psi_{b}\rangle$ from $|\sqrt{\pi}\rangle$. The ability to find a good point to short-jump to (i.e., constant $b$ where a jump takes $\operatorname{poly}(n)$ time to make) is where the inherent speedup over Grover comes from. If we just decided to do only the short-jump and sample until we found the ground state, the algorithm would only be quadratically slower due to amplitude amplification on the long-jump (assuming good gap costs $|\langle\psi_{b}|z^{*}\rangle|^{-1}$), than the full short-path algorithm.
The goal of this section is to determine conditions, using an initial Hamiltonian that is the discriminant of a reversible Markov chain, under which a short-jump can be done efficiently. However, it does not determine whether such a short-jump provides super-Grover runtime, which is the goal of the next subsection. The runtime of a short jump is captured by the short-path condition from [DPCB23, Has18a], where we present a natural generalization.
Definition 4.11 ($\theta$ Short-path condition).
Let $\Pi_{\perp}:=\mathbb{I}-|\sqrt{\pi}\rangle\!\langle\sqrt{\pi}|$. Then, the $\theta$ short-path condition holds for some constant $b>0$ if
$\textup{{GSE}}(\Pi_{\perp}H_{b}\Pi_{\perp})\geq-1+\theta$
where GSE denotes “ground-state energy”.
The short-path condition was used in previous papers to prove a variety of other sufficient conditions for super-Grover runtime. For example, the short path condition implies a lower bound on the spectral gap of $H_{b}$.
Lemma 4.12 ($\theta$ Short-path $\implies$ $\theta$ Spectral Gap Bound, adapted from Proposition 5 of [DPCB23]).
If $H_{b}$ satisfies the $\theta$ short-path condition, then the spectral gap of $H_{b}$ is at least $\theta$, i.e., all excited states have energy at least $-1+\theta$.
However, we stress that the short-path condition’s main purpose is to show the existence of an efficient “short jump”, i.e. $|\sqrt{\pi}\rangle\rightarrow|\psi_{b}\rangle$. In fact an inverse-$\operatorname{poly}(n)$ short-path condition is effectively equivalent to an inverse-$\operatorname{poly}(n)$ spectral gap at $H_{b}$ and overlap between $|\psi_{b}\rangle$ and $|\sqrt{\pi}\rangle$. Thus one can view it as a convenient way of combining the two conditions. The other consequence derived from the short-path condition was of a more technical nature. We summarize the implications of the short-path condition on the short jump in the following result.
Theorem 4.13 (Sufficient conditions for Efficient Short Jump).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain with spectral gap that is at least inverse-polynomial in $n$, and $H$ is a cost Hamiltonian satisfying the $\theta$ short-path condition at $b$ independent of the problem size $n$. If $\theta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, then there exists quantum algorithm for preparing an $\varepsilon$-approximation to $|\psi_{b}\rangle$ starting with $|\sqrt{\pi}\rangle$, which makes $\textup{poly}(n,\log(1/\varepsilon))$ queries to block-encodings of $D(P)$ and $H_{b}$.
Proof.
The result will follow if we can show that
$\min(\text{Gap}(-D(P)),\text{Gap}(H_{b}))\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert$ (12)
is inverse-polynomial in $n$. The assumption on the Markov chain gap implies that $\text{Gap}(-D(P))$ is inverse-polynomial. Lemma 4.12 implies that $\text{Gap}(H_{b})=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$.
The overlap is implied by the following argument. Let $|\psi_{b}^{\perp}\rangle$ be the component of $\psi_{b}$ orthogonal to $|\sqrt{\pi}\rangle$. The short-path condition implies that
$\langle\psi_{b}|\Pi_{\perp}H_{b}\Pi_{\perp}|\psi_{b}\rangle\geq-1+\Omega\left(\frac{1}{\operatorname{poly}(n)}\right),$ (13)
and since $\langle\psi_{b}|H_{b}|\psi_{b}\rangle\leq-1$, we have
$\lvert\langle\psi_{b}|H_{b}|\psi_{b}\rangle-\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\rvert=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right).$ (14)
Suppose $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert=o\left(\frac{1}{\operatorname{poly}(n)}\right)$, then since $\lVert H_{b}\rVert_{2}\leq 2$
$\lvert\langle\psi_{b}|H_{b}|\psi_{b}\rangle-\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\rvert=o\left(\frac{1}{\operatorname{poly}(n)}\right),$ (15)
a contradiction. ∎
We will later show that an even stronger condition on the overlap can be established assuming the spectral density condition and in this case $\langle\psi_{b}|\sqrt{\pi}\rangle=1-o(1)$.
If the conditions of Theorem 4.13 are satisfied, then we only need show that $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}$ is exponentially smaller (say by $\pi(E^{*})^{c/2}$ for some constant $c$) than $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}$ to achieve a super-qudratic speedup over Markov chain search. This is the task of bounding the runtime of the long jump.
In this paper and [DPCB23], the long-jump runtime is upper bounded by lower bounding the easier-to-handle quantity $\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\!\langle\psi_{b}|z^{*}\rangle\rvert$ by constructing an approximation to $|\psi_{b}\rangle\!\langle\psi_{b}|$. This makes use of a technical condition that ensures the ground state energy of $H_{b}$, $E_{b}$, does not decrease too much from the ground state energy of $-D(P)$. While it can be shown to hold via the short-path condition, we give intuition that it is a significantly weaker condition.
The following results provide generalized conditions under which the existence of a log-Sobolev and/or Poincaré inequality enable the spectral density condition (a very weak condition) to imply a constant $b$ at which a $\theta$ short-path condition exists, where $\theta$ is either the log-Sobolev or spectral gap of the chain $\mathcal{M}$. For Poincaré case, the spectral density condition is not even needed, although the spectral gap must of $\mathcal{M}$ must be constant.
From a bird’s-eye view, the role of the functional inequality is to upper bound a metric or divergence between $\psi_{b}^{2}$, $\ell_{2}$ distribution of $|\psi_{b}\rangle$ in the computational basis, and $\pi$. The variational definition of the corresponding metric or divergence plays the role of lower bounding. Together, these bounds are sufficient to derive a range of $b$’s where short-path holds.
Lemma 4.14.
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain that satisfies a log-Sobolev inequality with constant $\omega$. Then, for all quantum states $|\psi\rangle$ one has
$\frac{1-\langle\psi|D(P)|\psi\rangle}{\omega}\geq\textup{KL}(\psi^{2}\|\pi),$
[…]
Proof.
From a straightforward calculation, it follows
$\mathcal{D}(\psi,\psi)$ $=\langle\psi,(I-P)\psi\rangle_{\pi}$ (16)
$=\mathbb{E}_{\pi}(\psi^{\mathsf{T}}\psi)-\mathbb{E}_{\pi}(\psi^{\mathsf{T}}P\psi)$ (17)
$=\sum_{x\in\mathcal{X}}(\sqrt{\pi(x)}\psi(x))^{2}-\sum_{x\in\mathcal{X}}\sqrt{\pi(x)}\psi(x)D(P)_{xy}\sqrt{\pi(y)}\psi(y)$ (18)
$=\lVert\sqrt{\pi}\psi\rVert_{2}^{2}-\langle\sqrt{\pi}\psi|D(P)|\sqrt{\pi}\psi\rangle.$ (19)
Now,
$\lVert\sqrt{\pi}\psi\rVert_{2}^{2}-\langle\sqrt{\pi}\psi|D(P)|\sqrt{\pi}\psi\rangle\geq\omega(\mathbb{E}_{\pi}(\psi^{2}\ln(\psi^{2}))-\mathbb{E}_{\pi}(\psi^{2}\ln(\mathbb{E}\psi^{2}))),$
and
$\mathbb{E}_{\pi}(\psi^{2}\ln(\psi^{2}))-\mathbb{E}_{\pi}(\psi^{2}\ln(\mathbb{E}\psi^{2}))=\sum_{x\in\mathcal{X}}\pi(x)\psi^{2}(x)\ln(\psi^{2}(x))-\lVert\sqrt{\pi}\psi\rVert_{2}^{2}\ln(\lVert\sqrt{\pi}\psi\rVert_{2}^{2}).$
Consider $\psi=\frac{\psi^{\prime}}{\sqrt{\pi}}$ with $\lVert\psi^{\prime}\rVert_{2}=1$. Then
$1-\langle\psi|D(P)|\psi\rangle\geq\omega\sum_{x\in\mathcal{X}}\psi^{2}(x)\ln\left(\frac{\psi(x)^{2}}{\pi(x)}\right)=\omega\text{KL}(\psi^{2}||\pi).$
Theorem 4.15 (Sufficient $b$ for spectral density to imply short path).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov Chain that satisfies an $\omega$ log-Sobolev inequality. If
$b<\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right)$
then $\gamma$ spectral density implies an $\frac{\omega}{2}$ short-path condition.
Proof.
Let $F(E)$ be the cumulative distribution function for the cost function $H$. Define $F_{\eta}(E)$ to be the cumulative distribution function for $g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$. Then,
$F_{\eta}(v)=\begin{cases}0&v<-1\\ F(E^{\ast}(1-\eta-\eta v))\leq(\pi(E^{\ast}))^{\gamma}&-1\leq v<0\\ 1&v\geq 0\end{cases}$
where we assume that the probability of low energy states under $\pi$ is low, i.e.,
$F((1-\eta)E^{\ast})=\mathbb{P}_{\pi}(E\leq(1-\eta)E^{\ast})\leq\left(\pi(E^{\ast})\right)^{\gamma}$
for $0<\gamma\leq 1$ where $\pi(E^{\ast})$ is given by
$\pi(E^{\ast}):=\mathbb{P}_{\pi}(E=E^{\ast})=\sum_{\begin{subarray}{c}\{x\in\mathcal{X}:E(x)=E^{\ast}\}\end{subarray}}\pi(x).$
If this bound does not hold, then for a very small $\eta$, there is a high probability mass for the states with energy closer to $E^{\ast}$. Therefore, we can find an approximate optimizer by randomly sampling from $\pi(x)$ in time sub-exponential in $\log(1/\pi(E^{\ast}))$. Note that when $\pi$ is uniform and the ground state is non-degenerate, we recover the original condition of [DPCB23, see, Lemma 5] since
$F((1-\eta)E^{\ast})\leq 2^{-\gamma n}.$
For a quantum state $|\psi\rangle$ let $\psi^{2}$ denote its $\ell_{2}$ distribution in the computational basis. Applying Donsker and Varadhan’s variational formula [DV83] for $\textup{KL}(\psi^{2}\|\pi)$, we may write
$\textup{KL}(\psi^{2}\|\pi)=\sup_{f}\{\mathbb{E}_{\psi^{2}}[f(x)]-\ln(\mathbb{E}_{\pi}[\exp(f(x))]\}.$
Choosing $f(x)=-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)$ and defining $U_{\psi}=\mathbb{E}_{\psi^{2}}[g_{\eta}(\frac{H}{E^{\ast}})]$, it follows
$\textup{KL}(\psi^{2}\|\pi)$ $\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)\mathbb{E}_{\psi^{2}2}\left[g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)\right]-\ln\left(\mathbb{E}_{\pi}\left[e^{-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}\left(\frac{H(x)}{E^{\ast}}\right)}\right]\right)$
$\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)U_{\psi}-1.$
The final inequality follows from
$\mathbb{E}_{\pi}\left[e^{-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}(\frac{H(x)}{E^{\ast}})}\right]$ $=\sum_{g_{\eta}(\frac{H(x)}{E^{\ast}}))=0}\pi(x)+\sum_{g_{\eta}(\frac{H(x)}{E^{\ast}})\neq 0}\pi(x)e^{-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)g_{\eta}(\frac{H(x)}{E^{\ast}})}$
$\leq\pi(E(x)=0)+\pi(E(x)\neq 0)e^{\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)}$
$=\pi(E(x)=0)+\left(\pi^{\ast}\right)^{\gamma}e^{\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)}$
$\leq 2,$
as we assume $\pi(g_{\eta}(\frac{H}{E^{\ast}})\neq 0)\leq(\pi^{\ast})^{\gamma}$ and $g_{\eta}(\frac{H}{E^{\ast}})\geq-1$.
Thus we have the following lower bound on the KL divergence:
$\textup{KL}(\psi^{2}\|\pi)\geq-\gamma\ln\left(\frac{1}{\pi^{\ast}}\right)U_{\psi}-1.$ (20)
We also have the following upper bound the KL divergence from Lemma 4.14:
$\frac{1-\langle\psi|D(P)|\psi\rangle}{\omega}\geq\textup{KL}(\psi^{2}\|\pi),$ (21)
where $\omega$ is the LS constant of $P$.
The following argument attempts to find an upper bound $b^{*}$ on the $b$, such that for all $b<b^{*}$ the two bounds above become contradicting if short-path is not satisfied.
Suppose for contradiction that the $\frac{\omega}{2}$ short-path condition is violated at $b$, i.e.
$\textup{{GSE}}(\Pi_{\perp}H_{b}\Pi_{\perp})<-1+\frac{\omega}{2},$
where $\omega$ is the log-Sobolev constant of $P$. If $|\psi_{b}^{\prime}\rangle$ is the ground state of $\Pi_{\perp}H_{b}\Pi_{\perp}$, then
$-1+\frac{\omega}{2}>\langle\psi_{b}^{\prime}|H_{b}|\psi_{b}^{\prime}\rangle=-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle+b\langle\psi_{b}^{\prime}|G_{\eta}|\psi_{b}^{\prime}\rangle=-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle+bU_{\psi_{b}^{\prime}},$ (22)
which implies
$1-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle<\frac{\omega-2bU_{\psi_{b}^{\prime}}}{2}.$ (23)
Thus if short-path is violated, then the KL upper bound above reduces to
$\frac{\omega-2bU_{\psi_{b}^{\prime}}}{2\omega}\geq\textup{KL}(\psi_{b}^{\prime 2}\|\pi).$
We also have generally that a Poincaré inequality (LS inequality implies Poincaré) implies that for any $|\psi\rangle$
$\langle\psi|-D(P)|\psi\rangle\geq-1+\delta\geq-1+\omega,$ (24)
where $\delta$ denotes the Poincaré constant of $P$, so if short-path is violated combining the above with Equation (22) gives
$\frac{\omega}{2b}<-U_{\psi_{b}^{\prime}},$ (25)
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $-U_{\psi_{b}^{\prime}}\leq 1$.
Hence, when the short path condition is violated, Equations (20) and (21) imply that
$0\geq-\gamma\ln\left(\frac{1}{\pi(E^{\ast})}\right)U_{\psi_{b}^{\prime}}-1-\frac{\omega-2bU_{\psi_{b}^{\prime}}}{2\omega}.$ (26)
As mentioned earlier, we want to a range of $b$’s that contradicts this inequality, and so we solve
$0<-\gamma\ln\left(\frac{1}{\pi(E^{\ast})}\right)U_{\psi_{b}}-1-\frac{\omega-2bU_{\psi_{b}^{\prime}}}{2\omega},$
which yields
$b<\frac{3\omega}{2U_{\psi_{b}^{\prime}}}+2\gamma\omega\ln(1/\pi(E^{\ast}))$
and using Equation (25) since we want to consider the smallest right hand side:
$b<\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right).$
Thus, the short path condition must hold for the values of $b$ given in theorem statement. ∎
It is reasonable to question if a simpler Poincaré inequality for $\mathcal{M}=(\mathcal{X},P,\pi)$ which only depends on the spectral gap $\delta$ of $P$ would suffice to get a bound on $b$. Unfortunately, it appears that this does not provide a useful bound unless $\delta$ is constant.
Lemma 4.16.
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a Markov chain that satisfies a Poincaré inequality with constant $\delta$. Then, for all quantum states $|\psi\rangle$ one has
$\frac{1-\langle\psi|D(P)|\psi\rangle}{\delta}\geq[\textup{TV}(\psi^{2},\pi)]^{2},$
where $\textup{TV}(\cdot,\cdot)$ denotes the total variation distance, and $\psi^{2}$ is the $\ell_{2}$ distribution of $|\psi\rangle$ in the computational basis.
Proof.
The proof roughly follows that of Lemma 4.14, i.e. using Equation (16), but instead uses $\pi$-Variance. From Poincaré:
$\frac{1-\langle\psi|D(P)|\psi\rangle}{\delta}=\frac{\mathcal{D}(\psi/\sqrt{\pi},\psi/\sqrt{\pi})}{\delta}$ $\geq\text{Var}_{\pi}[\psi/\sqrt{\pi}]=\mathbb{E}_{\pi}[\psi^{2}/\pi]-(\mathbb{E}_{\pi}[\psi/\sqrt{\pi}])^{2}$
$=1-|\langle\sqrt{\pi}|\psi\rangle|^{2}$
$\geq[\text{TV}(\pi,\psi^{2})]^{2}.$
Note the following does not make use of the spectral density condition, which is one reason for the weak upper bound on $b$. However, it does suffice for $P$ with constant spectral gaps.
Theorem 4.17 (Sufficient $b$ for short-path under Poincaré inequality).
Suppose $\mathcal{M}=(\mathcal{X},P,\pi)$ is a reversible Markov chain that satisfies a Poincaré inequality with constant $\delta$. If
$b<\delta\frac{4\sqrt{6}-1}{10},$
then, the $\frac{\delta}{2}$ short-path condition is satisfied.
Proof.
The proof follows a similar structure as Theorem 4.15. We have the variational definition of TV:
$\text{TV}(\pi,\psi)=\textup{sup}_{f:\lVert f\rVert\leq 1}\frac{1}{2}(\mathbb{E}_{\psi^{2}}[f(x)]-\mathbb{E}_{\pi}[f(x)]).$
We can choose $f(x)=-g_{\eta}(H(x)/E^{*})$, which satisfies $\|f\rVert\leq 1$, so
$\text{TV}(\pi,\psi^{2})\geq\frac{1}{2}(-U_{\psi}-1),$
since $\mathbb{E}_{\pi}[f(x)]\leq 1$. Note $U_{\psi}=\mathbb{E}_{\psi^{2}}[f(x)]$.
Thus, we have the following lower bound:
$[\text{TV}(\pi,\psi^{2})]^{2}\geq\frac{1}{4}(U_{\psi}+1)^{2}.$ (27)
We also have the following upper bound on the TV from Lemma 4.16:
$\frac{1-\langle\psi|D(P)|\psi\rangle}{\delta}\geq[\textup{TV}(\psi^{2}\|\pi)]^{2},$ (28)
where $\delta$ is the Poincaré constant of $P$.
Suppose for contradiction that the $\frac{\delta}{2}$ short-path condition is violated at $b$, i.e.
$\textup{{GSE}}(\Pi_{\perp}H_{b}\Pi_{\perp})<-1+\frac{\delta}{2},$
where $\omega$ is the log-Sobolev constant of $P$. Thus, if $|\psi_{b}^{\prime}\rangle$ is the ground state of $\Pi_{\perp}H_{b}\Pi_{\perp}$, then
$-1+\frac{\delta}{2}>\langle\psi_{b}^{\prime}|H_{b}|\psi_{b}^{\prime}\rangle=-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle+b\langle\psi_{b}^{\prime}|G_{\eta}|\psi_{b}^{\prime}\rangle=-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle+bU_{\psi_{b}^{\prime}},$ (29)
which implies
$1-\langle\psi_{b}^{\prime}|D(P)|\psi_{b}^{\prime}\rangle<\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2}.$ (30)
Thus if short-path is violated, then the KL upper bound above reduces to
$\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta}\geq[\textup{TV}(\psi_{b}^{\prime 2},\pi)]^{2}.$
If $P$ satisfies a $\delta$ Poincaré inequality, then we have that for any $|\psi\rangle$
$\langle\psi|-D(P)|\psi\rangle\geq-1+\delta,$ (31)
so if short-path is violated combining the above with Equation (29) gives
$\frac{\delta}{2b}<-U_{\psi_{b}^{\prime}},$ (32)
where $U_{\psi_{b}^{\prime}}<0$ by construction. Also, by construction $-U_{\psi_{b}^{\prime}}\leq 1$.
Hence, when the short path condition is violated, Equations (27) and (28) imply that
$0\geq\frac{1}{4}(U_{\psi_{b}^{\prime}}+1)^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta}.$
As mentioned earlier, we want to a range of $b$’s that contradicts this, so we solve
$0<\frac{1}{4}(U_{\psi_{b}^{\prime}}+1)^{2}-\frac{\delta-2bU_{\psi_{b}^{\prime}}}{2\delta},$
so
$0\leq\delta U_{\psi^{\prime}_{b}}^{2}+(2\delta+4b)U_{\psi^{\prime}_{b}}-3\delta$
and using Equation (32) since we want to consider the smallest r.h.s.:
$0\leq\delta^{2}-4\delta b-20b^{2},$
so using the positive root
$b<\delta\frac{4\sqrt{6}-1}{10}.$
Thus, the short-path condition must hold for the values of $b$ given in theorem statement. ∎
It is also natural to ask if a modified log-Sobolev inequality would work. The apparent issue is with relating the Dirichlet form to the energy with respect to $D(P)$. Although we could not prove a generic bound in terms of modified log-Sobolev constant, we can use it to derive a lower bound on standard Log Sobolev constant by using the following thoerem.
Theorem 4.18 (Theorem 1 in [Sal21]).
Let $\omega_{\textrm{LSI}}$ and $\omega_{\textrm{MLSI}}$ respectively denote the log-Sobolev constant and the modified log-Sobolev constant of a Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. If $\mathcal{M}$ is reversible, then
$\omega_{\textrm{LSI}}\geq\omega_{\textrm{MLSI}}/\log(1/p)$
where $p$ is the smallest non-zero element in $P$.
Note that for general Markov chains $p$ can be exponentially small. However, in certain cases such as single site Glauber dynamics on bounded degree graphs, $p$ is only polynomially small. Therefore, in these special cases, our main theorem implies a super quadratic speed up although the speedup term $c$ might be falling with $n$.
Properties of the Short Jump
We now show some properties of the ground state $|\psi_{b}\rangle$ that is obtained as a result of the short jump. In the next section, we will use an approximation ground state projector to bound the runtime of the long and short jumps together (as in  [DPCB23]). It is instructive however, to bound the runtime of the short jump alone and confirm that it indeed takes only polynomial time. In the following, we show that if we select $b$ such that a $\theta$-short path condition is satisfied, the time taken for the short jump is in fact $\mathcal{O}(\delta^{-1})$ where $\delta$ is the spectral gap of $P$. This analysis supports the discussion of the algorithm in Section 3. We first establish a technical condition on the change in ground state energy. An analogous bound is used in [DPCB23] to bound the runtime and while we will do the same, it is convenient to introduce it here as we establish further properties of $|\psi_{b}\rangle$.
Lemma 4.19 (Ground-state energy shift bound, adapted from Proposition 6 of [DPCB23]).
Suppose that $\gamma$-spectral density holds and the $\frac{\theta}{2}$-short-path condition holds at $b$, then
$\lvert E_{b}\rvert<1+\frac{4(\pi(E^{\ast}))^{\gamma}}{\theta}.$ (33)
The above lemma utilizes the following expression for $E_{b}$:
$E_{b}$ $=\langle\psi_{b}|H_{b}|\psi_{b}\rangle=-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-b^{2}\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle,$
where
$W_{b}$ $:=(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp})^{-1}.$
The main purpose of this lemma is for determining the cost of the long-jump, which we do in the next section. Specifically, we will want to bound the magnitude of the energy shift away from two. We provide intuition why doing so does not require a strong ($\theta$ being inverse-polynomial in $n$) short-path condition. In the above lemma, the short-path condition is used to bound the spectral norm of $W_{b}$, where a $\theta$ short-path condition implies $\lVert W_{b}\rVert_{2}\leq\theta^{-1}$. However, the actual quantity of interest is $\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle$, where $\lVert G_{\eta}|\sqrt{\pi}\rangle\rVert_{2}\leq\pi(E^{\ast})^{\gamma}$. For “hard” problems $\pi(E^{\ast})^{\gamma}$ will be exponentially small in the problem size, and so $\lVert W_{b}\rVert$ can be even exponentially large for a small ground-state energy shift. Thus, for such problems, the short-path condition we need for a good gap ($\theta$ being inverse-polynomial in $n$) is stronger than the short-path condition needed for a small-enough energy shift.
If we do have an inverse-polynomial short-path condition, then the above lemma places strong conditions on the quantum state obtained by the short jump, namely that its trace distance from the starting state is small whenever hard to solve exactly by sampling from $\pi$. Specifically, we have the following lemma
Lemma 4.20.
Suppose that the $\gamma$ spectral density and $\frac{\delta}{2}$-short path condition hold for some constant $b$. It holds that $\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\mathcal{O}\left(\delta^{-1}(\pi(E^{\ast}))^{\gamma/2}\right)$ where $\delta$ is the spectral gap of $P$. Consequently, if $\delta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right),\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\mathcal{O}^{\ast}\left((\pi(E^{\ast}))^{\gamma/2}\right)$.
Proof.
Note that the ground state energy of $H_{b}$ is $\leq-1$ since $H_{b}-(-D(P))$ is negative definite. It follows from Lemma 4.19 and the conditions of this lemma that $-1-\mathcal{O}^{\ast}(\pi(E^{\ast}))^{\gamma})\leq E_{b}\leq-1$. We observe that
$0\leq-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle-\langle\psi_{b}|H_{b}|\psi_{b}\rangle=\mathcal{O}(\delta^{-1}\pi(E^{\ast})^{\gamma}),$
$\implies$ $\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle=\mathcal{O}(\delta^{-1}\pi(E^{\ast})^{\gamma})+b\langle\psi_{b}|G_{\eta}|\psi_{b}\rangle$
$\implies$ $\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert=\mathcal{O}(\delta^{-1}(\pi(E^{\ast})^{\gamma}),$
where the last equality follows from the negative semi-definiteness of $G_{\eta}$.
We may write the state $|\psi_{b}\rangle$ as $\alpha|\sqrt{\pi}\rangle+\sqrt{1-\alpha^{2}}|\sqrt{\pi}^{\perp}\rangle$ where $|\pi^{\perp}\rangle$ is a quantum state such that $\langle\sqrt{\pi}^{\perp}|\sqrt{\pi}\rangle=0$ (We may take $\alpha$ to be real and in the interval $[0,1]$ since $|\psi_{b}\rangle,|\sqrt{\pi}\rangle$ are the ground states of stoquastic Hermitian matrices). It follows from the definition of trace distance that $\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\sqrt{1-\alpha^{2}}$. From the definition of spectral gap, and observing that ${\sqrt{\pi}}$ is the ground state of $-D(P)$, we have
$\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert$
$=$ $\lvert(1-\alpha^{2})\left(\langle\sqrt{\pi}^{\perp}|D(P)|\sqrt{\pi}^{\perp}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\right)\rvert\geq\delta(1-\alpha^{2}).$
From our two bounds on $\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert$, it follows that
$\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\sqrt{1-\alpha^{2}}=\mathcal{O}\left(\delta^{-1}(\pi(E^{\ast}))^{\gamma/2}\right)=\mathcal{O}^{\ast}\left(\pi(E^{\ast}))^{\gamma/2}\right),$
which completes the proof. ∎
Since $|\sqrt{\pi}\rangle$ and $|\psi_{b}\rangle$ are pure states, we obtain a bound on the overlap between the states as an immediate consequence of the above.
Corollary 4.21.
Suppose that the $\gamma$ spectral density and $\frac{\delta}{2}$-short path condition hold for some constant $b$ with $\delta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, where $\delta$ is the spectral gap of $P$. Then, $|\langle\psi_{b}|\sqrt{\pi}\rangle|=1-\mathcal{O}^{\ast}((\pi(E^{\ast})^{\gamma})$.
It is also clear from the above that if $\pi(E^{\ast})^{-1}$ is super-polynomial in $n$ (as is the case for the problems considered here), the trace distance between $|\psi\rangle$ and $|\psi_{b}\rangle$ decays super-polynomially in $n$. From the operational definition of trace distance, the difference in probability of obtaining a specific outcome when performing a fixed measurement on two states $\rho_{1},\rho_{2}$ is upper bounded by their trace distance. Since the trace distance is monotonic under discarding of identical subsystems, it also holds for all $m\in\mathbb{N}$ that $\lVert\rho_{1}^{\otimes m}-\rho_{2}^{\otimes m}\rVert_{\mathrm{Tr}}\leq\lVert\rho_{1}-\rho_{2}\rVert_{\mathrm{Tr}}$. As a consequence of the above considerations, and the monotonicity of trace distance we have the following corollary, which illustrates that the short jump by itself does not offer any advantage if used in a polynomial time approximation algorithm.
Corollary 4.22.
Suppose that the conditions of Lemma 4.20 are satisfied, and let $\mathcal{A}$ be any non-adaptive quantum or classical algorithm applied to $\operatorname{poly}(n)$ copies of an input quantum state. Each outcome of $\mathcal{A}$ that is observed with probability $p_{1}$ when the algorithm is applied to $|\psi_{b}\rangle$, must be observed with probability $p_{2}=p_{1}\pm\mathcal{O}^{\ast}((\pi(E^{\ast})^{\gamma})$.
4.2.2 The Long Jump
The long jump is the preparation of an optimal solution in $\Pi^{\ast}$ from $|\psi_{b}\rangle$, where ideally $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$. As discussed in the previous subsection, the only quantum speedup from this step is quadratic and is due to amplitude amplification. That is, the quantum short-jump plus classical sampling would cost $\mathcal{O}^{\ast}(\pi(E^{*})^{-(1-c)})$, provided that the overlap condition just mentioned is satisfied. Accordingly, a quantum short-jump applied to the discriminant of the Markov Chain in Markov Chain search could still provide a nontrivial speedup. The goal of this subsection is to determine conditions on the cost Hamiltonian in terms of the Markov Chain $\mathcal{M}$ under which $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$ holds for some constant $c$. If we combine with amplitude amplification we get the runtime stated for the generalized short-path algorithm.
In order to bound the terms in the runtime involving the ground state projector $|\psi_{b}\rangle\langle\psi_{b}|$, we follow [DPCB23] in approximating $|\psi_{b}\rangle\!\langle\psi_{b}|$ through the use of a simple $-\ell$ polynomial $\mathcal{P}_{\ell}$. This quantity is related to the runtime via
$\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert^{-1}+\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}\leq 2\lVert\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|\Pi^{\ast}\rVert_{2}^{-1}.$ (34)
Let $E_{b}$ denote the ground state energy of $H_{b}$.
Lemma 4.23 (Runtime bound by approximate projector).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain with $|\mathcal{X}|=V$. Define
$\mathcal{P}_{\ell}:=\left(\frac{H_{b}}{\lvert E_{b}\rvert}\right)^{\ell},$
and let $\omega$ be the log-Sobolev constant of $D(P)$. For either all even or all odd $\ell$ we have
$\left\langle\sqrt{\pi}\middle|\mathcal{P}_{\ell}\middle|z\right\rangle-V\left(1-\frac{\omega}{2}\right)^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z\right\rangle,$
for any assignment $z$. Note the same also holds with $\omega$ replaced by the spectral gap $\delta$.
We now recall the definition of $\alpha$-subdepolarizing from [DPCB23] but generalized to arbitrary Markov Chains.
Definition 4.24 ($\alpha_{P}$-subdepolarizing).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The pair $(\mathcal{M},H)$ satisfies the $\alpha$-subdepolarizing property if for $f(x)=-g_{\eta}(-x)$ (as defined earlier), the following holds for any set of constants $0<c_{1},\dots,c_{T}<1,\forall T\in\mathbb{N}$:
[…]
where $E^{\ast}$ is the ground state energy of $H$.
As shown in the proof of Proposition 3 of [DPCB23], if $\Delta_{P}(\eta)$-stability holds, then $\alpha_{P}$-subdepolarizing is satisfied. However, the converse also holds.
Lemma 4.25 ($\Delta_{P}(\eta)$ stable $\iff\alpha_{P}$ subdepolarizing ).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The pair $(\mathcal{M},H)$ satisfies the $\Delta_{P}(\eta)$-stable if and only if it satisfies $\alpha_{P}$-subdepolarizing. They are related by the equation $\alpha_{P}=\frac{\Delta_{P}(\eta)}{\lvert E^{\ast}\rvert(1-\eta)}$.
We now briefly remark on some useful upper bounds on $\Delta_{P}(\eta)$ that we alluded to earlier. However, using too loose of an upper bound may result in the runtime analysis not indicating a speedup. Note that choice of the bound on $\Delta_{P}(\eta)$ is not actually used by Algorithm 2.
Lemma 4.26 (Upper bounds on $\Delta(\eta)$).
If $H$ has $P$ pseudo-Lipschitz norm $\lVert H\rVert_{P}$, then for $\eta\in[0,1)$,
$\sqrt{\lVert H\rVert_{P}}\geq\Delta_{P}(\eta).$
Furthermore, if
[…]
then for $\eta\in[0,1)$
$\tilde{\Delta}_{P}\geq\Delta_{P}(\eta).$
Our next result generalizes [DPCB23, Lemma 3], which uses $\alpha_{P}$ subdepolarizing (or equivalently $\Delta_{P}(\eta)$) to lower bound $\lvert\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z^{\ast}\rangle|$.
Lemma 4.27 (Overlap with general Markov chain).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. Given positive parameters $\eta<1$, $b<1$, $\alpha<(1-b)/2$, and integer $\ell$, suppose that $(H,g_{\eta})$ has the $\alpha$-subdepolarizing property, $3/\alpha^{2}\leq\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, and that $H_{b}$ satisfies the small-ground-energy shift condition. Define the function $F:[0,1]\mapsto[0,1]$ as $F(x):=1-x+x\ln(x)$. Let $z^{\ast}\in\{-1,+1\}^{n}$ be an optimal assignment, i.e. $H(z^{\ast})=E^{\ast}$. Then,
$\braket{\pi}{\mathcal{P}_{\ell}}{z^{\ast}}\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2}).$
Proof.
Define $A,B$, and $f$ by the following equations:
$A$ $=D(P)$
$B$ $=-bg_{\eta}\left(\frac{H}{E^{\ast}}\right)=bf\left(\frac{H}{E^{\ast}}\right),$
so
$\langle y|A|x\rangle=\pi^{1/2}(x)P(x,y)\pi^{-1/2}(y).$
Then, the approximate projector can be written as
$\mathcal{P}_{\ell}=\frac{(-A-B)^{\ell}}{|E_{b}|^{\ell}}.$
By Lemma 33, if $\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, then we can take $|E_{b}|^{\ell}=\Theta(1)$ and ignore the denominator. To compute the numerator, start with
$\braket{\pi}{A}{z^{\ast}}=\braket{\pi}{z^{\ast}}=\pi(z^{\ast})^{1/2}$
$\braket{\pi}{B}{z^{\ast}}=b\pi(z^{\ast})^{1/2}.$
Next, we compute
[…]
$=b\sum_{x_{1}\dots x_{k}}\braket{\pi}{B}{x_{k}}\pi(x_{k})^{-1/2}P(x_{k-1},x_{k})\pi(x_{k-1})^{1/2}\cdots\pi(x_{1})^{-1/2}P(z^{\ast},x_{1})\pi(z^{\ast})^{1/2}$
$=b\sum_{x_{1}\dots x_{k}}f\left(\frac{H(x_{k})}{E^{\ast}}\right)P(z^{\ast},x_{1})\cdots P(x_{k-1},x_{k})\pi(z^{\ast})^{1/2}$
$=b\pi(z^{\ast})^{1/2}\mathbb{E}_{x_{1}\sim z}\cdots\mathbb{E}_{x_{k}\sim x_{k-1}}f\left(\frac{H(x_{k})}{E^{\ast}}\right)$
$\geq b\pi(z^{\ast})^{1/2}f\left((1-\alpha)^{k}\right).$
In general, we can write any string of $A$’s and $B$’s as
$\dots AB^{c_{3}}AB^{c_{2}}AB^{c_{1}}AB^{c_{0}},$
for $c\in\ell_{1}(\mathbb{N}^{\infty})$, i.e., finite sequences of natural numbers of unbounded length. Let $\tilde{f}(x):=f(H(x)/E^{\ast}).$ Accordingly, we can compute
[…]
$=\sum_{x_{1},\dots}b^{\sum_{j}{x_{j}}}\dots\braket{x_{2}}{D}{x_{1}}f(H(x_{1})/E^{\ast})^{c_{1}}\braket{x_{1}}{D}{z^{\ast}}$
$=\pi(z^{\ast})^{1/2}b^{\sum_{j=0}^{\infty}{c_{j}}}\sum_{x_{1},\dots}\dots P(x_{1},x_{2})f(H(x_{1})/E^{\ast})^{c_{1}}P(x_{1},z^{\ast})$
[…]
$\geq\pi(z^{\ast})^{1/2}b^{\sum_{j=0}^{\infty}c_{j}}\prod_{j=0}^{\infty}f((1-\alpha)^{j})^{c_{j}}.$
By assumption, $b^{\sum_{j=0}^{\infty}c_{j}}$ is finite.
We only pick up $\pi(z^{\ast})^{1/2}$ term instead of $2^{-n/2}$ in front of the product. Therefore Propositions 15 and 16 of [DPCB23] hold. Hence, we obtain the stated result. ∎
The following uses the above results to bound the complexity of the short and long jumps. As mentioned earlier, the runtimes of the two jumps are bounded together due to the ease of analysis.
Theorem 4.28.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain, and suppose $|\mathcal{X}|=V$. If
$\ell~{}\geq~{}\max\left(\frac{3}{\alpha^{2}},\max_{z^{\ast}}4\frac{\ln(V/\sqrt{\pi(z^{\ast})})}{\omega}\right)$
and $\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, then
$\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert^{-1}+\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}=\mathcal{O}\left([\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{b\eta}{2\ln(1/\pi(E^{\ast}))\alpha}\right)}\right).$
A direct consequence is that the overall complexity in terms of queries to block-encodings of $D(P)$ and $H$ is
$\mathcal{O}\left([\min(\textup{Gap}(D(P)),\textup{Gap}(H_{b}))]^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{b\eta}{2\ln(1/\pi(E^{\ast}))\alpha}\right)}\right).$
Proof.
By Lemmas 4.23 and 4.19, for any optimal assignment $z^{\ast}$ we have
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2})-Ve^{-\omega\ell/2}.$
Consider $\ell\geq\max_{z^{\ast}}4\frac{\ln(V/\sqrt{\pi(z^{\ast}}))}{\omega}$, where the max is over optimal assignments, then
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2})-\pi(z^{\ast}).$ (35)
Note that $\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)>1$, so the first term on the right hand side will dominate asymptotically. In fact $\frac{F(1-\eta)}{\eta}\geq\eta/2$, so we have
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\Omega\left(\sqrt{\pi(z^{\ast})\exp\left(\frac{b\eta}{\alpha}\right)}\right).$
This clearly gives that
$\lVert\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|\Pi^{\ast}\rVert_{2}^{-1}\geq\Omega\left(\sqrt{\pi(E^{\ast})\exp\left(\frac{b\eta}{\alpha}\right)}\right).$
The result follows by using Equation (34). ∎
It may not be immediately obvious that the conditions on $\ell$ are not contradicting. Here, we give intuition for why this is not the case for the typical applications of the algorithm. Note for an efficient algorithm, at the very least we will need $\omega^{-1}=\mathcal{O}(\operatorname{poly}(n))$. For hard problems, $\pi(E^{\ast})$ will be exponentially small in $n$, thus the $\mathcal{O}\left(\min_{z^{\ast}}\frac{\omega}{(\pi(z^{\ast}))^{\gamma}}\right)$ upper bound on $\ell$ will be significantly larger than the lower bound. Note $1/\alpha^{2}$ is significantly smaller than $1/\pi(E^{\ast})$ for a hard problem, for showing super-Grover runtime we will want $\alpha=\Theta(\frac{1}{\ln(1/\pi(E^{\ast}))})$ anyways. Thus it is fine to assume we have the conditions on $\ell$ stated in Theorem 4.28 in settings where the algorithm can successfully be applied.
As should be apparent from Theorem 4.3, at a constant $b$ the existence of a speedup over Grover is determined solely by $\frac{\Delta_{P}}{\lvert E^{*}\rvert}$. Clearly, we must have that $\frac{\Delta_{P}}{\lvert E^{*}\rvert}=\Omega\left(\frac{1}{\ln(1/\pi(E^{*}))}\right)$, at least for a problem with at least exponential runtime. If this does not hold, then the runtime goes to zero asymptotically, an absurdity. However, it has not been directly shown that the derived runtime cannot lead to this contradiction. To put one’s mind at ease, we present the following result.
Lemma 4.29.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. Suppose $b^{*}$ and $\gamma$ in Theorem 4.3 are constants, and $\lVert H\rVert_{2}=\lvert E^{*}\rvert$. If
[…]
then $\frac{\tilde{\Delta}_{P}}{\lvert E^{*}\rvert}=\Omega\left(\frac{1}{\ln(1/\pi(E^{*}))}\right)$.
Proof.
Let $n$ be a real parameter that parameterizes the space of feasible states $\mathcal{S}(n)$ with size $|\mathcal{S}(n)|=S(n)$. Let $M(n)$ denote the mixing time of a Markov chain $P$ with transition density $\pi$ on $\mathcal{S}(n)$ such that for any $t\geq M(n)$, $\operatorname{TV}(P^{t}\delta_{x},\pi)\leq\frac{1}{100}$ for any $x\in\mathcal{S}(n)$.
Assume that $\Delta$ is a global upper bound, such that for all $x\in\mathcal{S}(n)$, $\mathbb{E}_{y\sim_{P}x}[H(y)]\leq H(x)+\Delta$. It is easy to observe that for any random variable $X$ taking values in $\mathcal{S}(n)$, it holds that $\mathbb{E}_{Y\sim_{P}X}[H(Y)]\leq\mathbb{E}_{X}[H(X)]+\Delta$. Now let $x_{\ast}$ be a global minimum of $H$ (with corresponding energy $E^{\ast}$) and consider taking $T=\lceil M(n)\rceil$ steps of $P$ starting from $x^{\ast}$, with the random state after $t\in[1,T]$ steps being denoted $X_{t}$. By induction, it is easy to see that $E_{X_{T}}[H(X_{T})]\leq E^{\ast}+\Delta T$. From the definition of mixing time however, it follows that $\mathbb{E}_{X_{T}}[H(X_{T})]\geq\mathbb{E}_{\pi}[H(x)]-\frac{\lvert E^{\ast}\rvert}{100}$. Denoting $\mathbb{E}_{\pi}[H(x)]$ by $\bar{E}$ it follows that $\Delta\geq\frac{\bar{E}-E^{\ast}}{T}-\frac{\lvert E^{*}\rvert}{100T}=\Omega\left(\frac{\lvert E^{*}\rvert}{M(n)}\right)$. So $\frac{\tilde{\Delta}_{P}}{\lvert E^{\ast}\rvert}=\Omega\left(\frac{1}{M(n)}\right)$.
If $b^{*}$ and $\gamma$ are constant, then the log-Sobolev constant must satisfy $\omega=\Omega\left(\frac{1}{\ln(1/\pi(E^{*})}\right)$. It follows from standard results on Markov Chains that $\omega\leq\frac{1}{M(n)}$. ∎
Note that in [DPCB23], the authors state an additional technical condition that is of course easy to satisfy, which is $\mathbb{E}_{\pi}[H(x)]=0$. The main reason for setting $\mathbb{E}_{\pi}[H(x)]=0$ is to ensure that $E^{\ast}<0$ and for the ease of proving the tail bounds. In our setting, for arbitrary $\pi$, the expectation may need to be estimated if used as a shift. However, the shift is not necessary to run the algorithm if the $E^{\ast}<0$ condition is already satisfied. The mean just appears as component of the runtime. Also, Corollaries 4.9 and 4.10 for the tail bounds do not assume this shift.
5 Applications of Generalized Short-Path Framework
5.1 Optimization with Fixed Hamming Weight: Transposition Walk
The $k$-particle Bernoulli-Laplace diffusion or Transposition Walk on $n$ sites is a random walk on the space of Hamming weight $k$ bistrings. For our case we will work with ±$\pm 1$ strings or “spin configurations” $x$ and define the Hamming weight $\lvert x\rvert$ as the number of $+1$’s. Formally, $\mathcal{X}=\{x\in\{-1,1\}^{n}:\lvert x\rvert=k\}$. A single step of BL consists of choosing, uniformly at random, a transposition that swaps some $x_{j}=1$ with another $x_{i}=-1$.
There is a very natural quantum Hamiltonian on $n$-qubits that encodes the discriminant of the transposition walk:
$D(P)=P=\frac{1}{k(n-k)}\sum_{i<j}\frac{X_{i}X_{j}+Y_{i}Y_{j}}{2},$
where $X_{j},Y_{j}$ denote the Pauli operators applied to qubit $j$. This is commonly called the complete-graph $XY$ mixer. Note that there is equality between $D(P)$ and $P$ because the walk is symmetric. The ground state of $-D(P)$ is the uniform superposition over Hamming-weight $k$ computational basis states, and thus encodes the stationary distribution of the transposition walk. Note that $|\sqrt{\pi}\rangle$ is just the Hamming-weight $k$ Dicke state, which can be prepared efficiently [BE19].
We have the following log-Sobolev inequality for the transposition walk.
Theorem 5.1 ([LY98, Sal21, Theorem 5], discrete-time).
Let $P$ be the transition matrix for $k$-particle Bernoulli-Laplace diffusion on $n$ sites and $\pi$ the stationary distribution. It then holds for any real valued function $\psi$ that
$\mathcal{D}(\psi,\psi)\geq\frac{n}{k(n-k)\tau_{LS}}\mathrm{Ent}(\psi^{2}).$
There is also a universal constant $\tau_{0}$ such that
$\tau_{LS}\leq\tau_{0}\log\left(\frac{n}{\min(k,n-k)}\right).$
This leads to the following bound on $b^{*}$ using the formula in Theorem 4.3.
Lemma 5.2.
For all $k$, we have the following bound on $b$ for the transposition mixer:
$b^{*}=\frac{2C\gamma}{3},$ (37)
for some constant $C$.
Proof.
We have
$\mathcal{D}(\psi,\psi)\geq\frac{n}{k(n-k)\tau_{0}\log\left(\frac{n}{\min(k,n-k)}\right)}\mathrm{Ent}(\psi^{2}).$ (38)
Thus
$b<\frac{2n\log_{2}\binom{n}{k}\gamma}{3k(n-k)\tau_{0}\log_{2}\left(\frac{n}{\min(k,n-k)}\right)}.$
Using that for $k=o(n)$, $\log_{2}\binom{n}{k}=\Theta(k\log_{2}(n/k))$, and $\frac{n}{\min(k,n-k)}=\Theta(\log_{2}(n/k))$, we get
$b<\frac{C2\gamma}{3},$ (39)
for some constant $C$ to be determined, so $b$ is constant for $k=o(n)$.
For $k=\Theta(n)$, we have that $\log_{2}\binom{n}{k}=\Theta(n)$, $\frac{n}{\min(k,n-k)}=\Theta(1)$, so $b^{*}$ is also constant. ∎
This leads to the following simple result that follows from applying Theorem 4.3.
Theorem 5.3.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}$. If $\lVert H\rVert_{P}=\mathcal{O}(1)$, and
$\lvert E^{\ast}\rvert=\Theta\left(\ln\binom{n}{k}\right),$
then there exists a short-path algorithm with runtime
$\mathcal{O}^{\ast}\left(\binom{n}{k}^{\left(\frac{1}{2}-\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\ln\binom{n}{k}\Delta_{P}}\right)}\right).$
Proof.
For all $k$ we have $\omega=\Omega\left(\frac{1}{k\ln(n/k)}\right)$, and $k(\ln(n/k))\ln(1/\pi(E^{*}))\asymp(\ln\binom{n}{k})^{2}$. Thus if $\lVert H\rVert_{P}=\mathcal{O}(1)$ and $\lvert E^{\ast}\rvert=\Theta\left(\ln\binom{n}{k}\right)$, then $\gamma$ is constant so then $b^{*}$ is. We also have that $\frac{\lvert E^{*}\rvert}{\Delta_{P}}=\mathcal{O}\left(\ln\binom{n}{k}\right)$, leading to the runtime presented. ∎
We apply the above result to a Hamming-weight constrained version of MaxCut over Erdős-Rényi graphs, which we call MaxCut-Hamming. One well-known special case is Hamming-weight $\frac{n}{2}$ called MaxBisection.
5.1.1 Hamming-weight Constrained MaxCut
Consider a graph $G(\mathcal{N},\mathcal{E})$ with vertex set $\mathcal{N}:=[n]$ and edge set $\mathcal{E}$. We assume $G$ is drawn from the Erdős-Rényi ensemble $\mathcal{G}\left(n,\frac{p}{n-1}\right)$ for a constant $p$, i.e., each edge $e_{ij}\in\mathcal{E}$ for $(i,j)\in[n]\times[n]$ is created with probability $\frac{p}{n-1}$ such that $G$ has an average degree of $p$. We are interested in solving the Maximum Bisection problem:
$\mathcal{C}^{*}_{\frac{n}{2}}:=\min_{x\in\{-1,1\}^{n}}\left\{-\frac{1}{2}\sum_{i<j}e_{ij}(1-x_{i}x_{j}):\lvert x\rvert=\frac{n}{2}\right\},$ (MaxBisection)
where $e_{ij}$ is a $\frac{p}{n-1}$ Bernoulli indicating whether the $(i,j)$ edge is present.
For generality, we strive to present the results for an arbitrary Hamming weight constraint of size $k$ and specify $k=\frac{n}{2}$ where necessary. We call the case where $k$ can be arbitrary the MaxCut Hamming problem:
$\mathcal{C}^{*}_{k}:=\min_{x\in\{-1,1\}^{n}}\left\{-\frac{1}{2}\sum_{i<j}e_{ij}(1-x_{i}x_{j}):\lvert x\rvert=k\right\}.$ (MaxCut-Hamming)
The following result shows that $\lVert H\rVert_{P}=\mathcal{O}(1)$.
Lemma 5.4.
For the MaxCut-Hamming Hamiltonian $H$, the pseudo Lipschitz constant $\lVert H\rVert_{P}$ under the transposiiton walk is $\mathcal{O}(1)$ with high probability over the graph.
The proof of the above lemma is deffered to the appendix. Next we show the existance of a tail bound for MaxBisection.
Lemma 5.5.
Let $\mathcal{X}=\{x\in\{-1,1\}^{n}:|x|=\frac{n}{2}\}$ and $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain. For (MaxBisection) on a graph $G\sim\mathcal{G}\left(n,\frac{p}{n-1}\right)$ and $D(P)$ being the transposition mixer, we have that
$\frac{4((1-\eta)((1-\eta)+\frac{p}{2})^{2}}{\tau_{0}}\lesssim\gamma.$
Proof.
Recall the expression for $\gamma$ in terms of the Herbst argument provided in Theorem 4.7:
$\gamma=\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert\psi\rVert_{P}\ln(1/\pi(E^{\ast}))}.$
From Lemma 5.4 we have that $\lVert H\rVert_{P}=\mathcal{O}(1)$. Applying Equation (38) for $k=\Theta(n)$, the log-Sobolev constant $\omega$ satisfies:
$\omega\geq\frac{n}{k(n-k)\tau_{0}\log\left(\frac{n}{\min(k,n-k)}\right)}=\frac{4}{\tau_{0}n}.$
From Lemma B.4 and Lemma B.1 we have
$((1-\eta)E^{\ast}-\mathbb{E}_{\pi}(H_{c}))^{2}\asymp\left[(1-\eta)n+\frac{np}{2}\right]^{2}.$
The definition of $\pi$ combined with the asymptotics of the binomial coefficient for $k=\Theta(n)$ gives
$\ln(1/\pi(E^{\ast}))\asymp n.$
Putting everything together:
$\frac{4((1-\eta)((1-\eta)+\frac{p}{2})^{2}}{\tau_{0}}$ $\lesssim\gamma.$
Lastly, $\lvert E^{\ast}\rvert=\Theta(n)$ with high probability from Lemma B.4. Since this is $\Theta\left(\ln\binom{n}{k}\right)$ for $k=\Theta(n)$ all of the conditions of Theorem 5.3 are met.
Unfortunately, the current analysis is insufficient to show this for $k=o(n)$. For example, using Lemma B.2 we can take an upper bound of
$\Delta_{P}=\frac{\mathcal{C}^{*}_{k}(n-2)}{k(n-k)},$ (40)
which gives that $\frac{\lvert E^{*}\rvert}{\Delta_{P}}=\mathcal{O}(k)$. However, $\ln\binom{n}{k}=\Theta(k\ln(n/k))$ for $k=o(n)$. Assuming a tail bound, this leads to a speedup that is falling with $n$. Specifically, the speedup is falling with $\frac{1}{\ln(n)}$, which is reminiscent of the running time achieved by [Has18a]. We summarize the two cases in the following theorem
Theorem 5.6.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian encoding the MaxCut-Hamming cost function for a graph $G\sim\mathcal{G}\left(n,\frac{p}{n-1}\right)$. Then either of the following runtimes hold depending on $k\leq n$.
  • If $k=\Theta(n)$, then there exists a short-path algorithm with runtime
    $\mathcal{O}^{\ast}\left(\binom{n}{k}^{\frac{1}{2}-c}\right),$
    for some constant $c$, and
  • if $k=o(n)$, then there exists a short-path algorithm with runtime
    $\mathcal{O}^{\ast}\left(\binom{n}{k}^{\frac{1}{2}-\frac{c}{\ln(n)}}\right),$
    for some constant $c$.
5.2 Glauber Dynamics
Glauber dynamics is a Markov Chain algorithm designed to sample from the Gibbs distribution of a system, particularly in spin models like the Ising or hard-core model [Gla63]. The Gibbs measure $\pi$ for a system with configuration space $\mathcal{X}$ and Hamiltonian $H(x)$ is defined as
$\pi(x)=\frac{\exp(-\beta H(x))}{Z(\beta)}$
where $x\in\mathcal{X}$ is a configuration, $\beta=\frac{1}{T}$ is the inverse temperature, and $Z(\beta)$ is the partition function
$Z(\beta)=\sum_{x\in\mathcal{X}}\exp(-\beta H(x)).$
Glauber dynamics generates a Markov chain with this measure as its stationary distribution by sequentially updating a single site (spin or vertex) according marginal distribution. Specifically, the update proceeds as follows: (i) a site (vertex) $v$ is chosen uniformly at random; (ii) the states of all other sites $u\neq v$ remain unchanged; (iii) the new state of $v$ is sampled from the marginal distribution of $v$ conditioned on its neighbors.
The efficiency of Gibbs sampling in this case is related to how fast the Glauber dynamics mixes to its stationary distribution. In fact, approximate Gibbs sampling is tightly connected to partition function estimation in terms of computational complexity [ŠVV09], both of which are closely related to statistical phase transitions. These transitions correspond to the uniqueness/non-uniqueness threshold on an infinite $d$-regular tree which captures whether the root of the tree is affected by the leaves. In the uniqueness regime, correlations decay rapidly, allowing efficient approximation of the partition function. However, beyond the non-uniqueness threshold, long range correlations emerge and no polynomial-time algorithm can approximate the partition function.
In particular, for the hardcore model with fugacity parameter $\lambda=e^{\beta}$, [Wei06] presented a fully polynomial time approximation scheme ($\mathsf{FPTAS}$) for computing the partition function a graphs with maximum degree $d$ when $\lambda\leq(1-\delta)\lambda_{c}$ where $\lambda_{c}=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$ is the corresponding uniqueness threshold. On the other hand,[Sly10] proved that is no fully polynomial randomized approximation scheme ($\mathsf{FPRAS}$) to approximate the partition function when $\lambda>\lambda_{c}$ unless $\mathsf{NP}=\mathsf{RP}$ confirming the main conjecture of [MWW09]. Similarly, for Ising model the phase transition occurs at $\beta_{c}=\frac{d-2}{d}$ for the antiferromagnetic case and $\beta_{c}=\frac{d}{d-2}$ for ferromagnetic case.
We first present our main result for the Glauber dynamics and then show that the necessary conditions hold for hardcore-model (maximum independent set problem) and Ising model.
Theorem 5.7.
Let $P$ be a Glauber dynamics chain on $\mathcal{X}$ that satisfies an $\omega$ log-Sobolev inequality with stationary distribution $\pi$, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
$E^{\ast}\coloneqq\min_{x\in\mathcal{X}}H(x).$
If $\|H\|_{P}=\mathcal{O}(1)$ and $\omega^{-1},\lvert E^{*}\rvert,$ and $\ln(1/\pi(E^{*}))$ are all $\Theta(n)$, then the short-path algorithm applied to the Glauber chain has a super-quadratic speedup over Markov Chain search with Glauber dynamics.
Proof.
This directly follows from application of Theorem 4.3. ∎
For convenience, we prove the following lemma that is useful when establishing $\|H\|_{P}=\mathcal{O}(1)$.
Lemma 5.8.
Let $P$ be a Glauber dynamics chain on $\mathcal{X}$ and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
$E^{\ast}\coloneqq\min_{x\in\mathcal{X}}H(x).$
If the following holds for all $x,x^{\prime}$ such that $P(x,x^{\prime})>0$
$\lvert H(x^{\prime})-H(x)\rvert=\mathcal{O}(1),$ (41)
then $\|H\|_{P}=\mathcal{O}(1)$.
Proof.
The proof simply follows from the definition of $\|H\|_{P}$ as
$\|H\|_{P}$ $=\max_{x}\sum_{x^{\prime}}P(x,x^{\prime})(H(x^{\prime})-H(x))^{2}\leq\max_{x,x^{\prime}}\lvert H(x^{\prime})-H(x)\rvert^{2}=\mathcal{O}(1).$
5.2.1 Maximum Independent Set Problem
Given a graph $G(\mathcal{N},\mathcal{E})$, an independent set is a subset of vertices where no two vertices are connected by an edge. A maximal independent set (MIS) is the largest independent set of $G$. Namely, we solve the optimization problem
$\mathcal{C}^{*}_{G=(\mathcal{N},\mathcal{E})}:=\min_{x\in\{0,1\}^{|\mathcal{N}|}}\left\{-\sum_{i\in\mathcal{N}}x_{i}:x_{i}+x_{j}\leq 1~{}\forall(i,j)\in\mathcal{E}\right\}.$ (Maximum Independent Set)
Equivalently, for $|\mathcal{N}|=n$, we can denote an independent set by a configuration $x\in\mathcal{X}\subseteq\{0,1\}^{n}$ such that if an edge $e=(i,j)\in\mathcal{E}$ then, $x_{i}x_{j}=0$. We denote the size of the maximum independent set by $|x^{\ast}|$. Then, finding the maximum independent set is equivalent to finding a ground state of Hamiltonian $H:\mathcal{X}\to[1,n]$,
$H(x)=-\sum_{i=1}^{n}x_{i}.$
The Hamiltonian $H$ is defined on constrained space $\mathcal{X}$ and therefore we need a constrained walk to explore the state space. For this purpose, we use the Glauber dynamics defined as
$P_{\lambda}(x,x^{\prime})=\begin{cases}0&\text{if }|x-x^{\prime}|>1\\ \frac{1}{n}\frac{\lambda}{\lambda+1}&\text{if }|x-x^{\prime}|=1\text{ and }x\subseteq x^{\prime}\\ \frac{1}{n}\frac{1}{\lambda+1}&\text{if }|x-x^{\prime}|=1\text{ and }x^{\prime}\subseteq x\\ 1-\sum_{x^{\prime\prime}\neq x}P_{\lambda}(x,x^{\prime\prime})&\text{if }x=x^{\prime}.\end{cases}$ (42)
This model is also referred as hard-core model in statistical physics and we’ll use the same terminology. Note that Glauber dynamics initialized at an independent set can only move between independent sets. Furthermore, it converges to its stationary distribution
$\pi_{\lambda}(x)=\frac{\lambda^{|x|}}{Z}.$
The parameter $\lambda$ is also called fugacity and one can recover the Gibbs form in 5.2 by setting $e^{\beta}=\lambda$. In principle, one can find the maximum independent set by setting $\lambda$ sufficiently high so that distribution $\pi_{\lambda}$ concentrates around the global minimum of $H$. Unfortunately, this approach results in exponential mixing time due to uniqueness/non-uniqueness phase transitions. Alternatively, one can draw exponentially many samples from $\pi_{\lambda}$ at $\lambda<\lambda_{c}$ by running the Glauber dynamics chain in polynomial time as Glauber dynamics mix efficiently below the critical threshold [CLV21]. As the current quantum techniques can only quadratically improve the run time of the first approach, we use the second apprach. More specifically, we consider Glauber chain $P_{\lambda}$ at $\lambda<\lambda_{c}$ so that we can prepare |$\ket{\pi_{\lambda}}$ efficiently. Next, we consider the short-path Hamiltonian $H_{b}:\mathcal{X}\mapsto\mathbb{R}$,
$H_{b}=-D(P_{\lambda})+bg_{\eta}\left(\frac{H}{|E^{\ast}|}\right),$
where $D$ is the discriminant matrix as usual and $E^{\ast}=\min_{x}H(x)=-|x^{\ast}|$. In accordance with the generalized short path framework, the algorithm starts from |$\ket{\pi_{\lambda}}$ and jumps to the ground state of $H_{b}$ for $b>0$. We describe how we can prepare the block-encoding of $D$ for Glauber dynamics and also prepare $\pi_{\lambda}$ in Appendix D.
The following lemma establishes the condition on $\|H\|_{P}$ given in theorem 5.7.
Lemma 5.9.
Let $P$ be a Glauber dynamics chain with fugacity parameter $\lambda$ on a graph $G(\mathcal{N},\mathcal{E})$. Then, $\|H\|_{P}=\mathcal{O}(1)$.
Proof.
As Glauber dynamics flips one spin at a time, for all $x,x^{\prime}\in\mathcal{X}$ such that $P(x,x^{\prime})>0$ we have
$|H(x)-H(x^{\prime})|\leq 1.$
Then, by Lemma 5.8, $\|H\|_{P}=\mathcal{O}(1)$. ∎
As required by Theorem 5.7, we need to characterize the log-Sobolev constant of Glauber dynamics for hard-core model on a graphs with a degree upper bounded by $d$. We start with the following fact from [CLV21],
Fact 5.10.
Let $V$ be a set of size $n$ and µ be a distribution over $[q]^{V}$. If $\pi$ satisfies the approximate tensorization of entropy with constant $C_{1}$ and $\pi$ is $u$-marginally bounded, then the Glauber dynamics for $\pi$ satisfies the standard log-Sobolev inequality with constant $\omega=\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, or $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
Proof.
Fix a configuration $x$ and consider a Markov chain $P_{v}$ that updates vertex $v$ according to marginal probability distribution $\pi_{v}=\pi(v|x_{V-\{v\}})$. Then, LS constant of this Markov chain $\rho_{v}$ is lower bounded by $\rho_{v}\geq\frac{1-2\pi_{v}^{\ast}}{\log(1/\pi_{v}^{\ast}-1)}$ when $\pi_{v}^{\ast}<\frac{1}{2}$ or $\rho_{v}=\frac{1}{2}$ when $\pi_{v}^{\ast}=\frac{1}{2}$ due to [DSC96, Theorem A.1]. By the definition of the log-Sobolev constant, we have
$\rho_{v}\operatorname{Ent}_{\pi_{v}}[f]$ $\leq\mathcal{E}_{P_{v}}(\sqrt{f},\sqrt{f})=\mathrm{Var}_{\pi_{v}}[\sqrt{f}].$
The transition matrix of Glauber dynamics can be written as $\frac{1}{n}\sum_{v\in V}\pi(v|\cdot)$. Therefore, the Dirichlet form for Glauber dynamics is
$\mathcal{D}(\sqrt{f},\sqrt{f})=\frac{1}{n}\sum_{v\in V}\mathrm{Var}_{\pi_{v}}[\sqrt{f}].$
Using the tensorization of entropy,
$\operatorname{Ent}_{\pi}[f]$ $\leq C_{1}\sum_{v\in V}\operatorname{Ent}_{\pi_{v}}[f]\leq C_{1}\sum_{v\in V}\frac{1}{\rho_{v}}\mathrm{Var}_{\pi_{v}}[\sqrt{f}]\leq C_{1}n\max_{v}\left({\frac{1}{\rho_{v}}}\right)\mathcal{D}(\sqrt{f},\sqrt{f}).$
Combining the marginal boundedness property of $\pi$ (i.e., $\pi_{v}\geq u$ for all $v\in V$), with the monotonicity of the function $\frac{1-2y}{\log(1/y-1)}$ for $y\in[0,1/2]$, it follows that $\omega\geq\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, and $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
Theorem 5.11 (Entropy factorization, Theorem 2.9 in [CLV21]).
Let $d\geq 3$ be an integer and $b,\eta>0$ be reals. Suppose that $G=(\mathcal{N},\mathcal{E})$ is an n-vertex graph of maximum degree at most $d$ and $\mu$ is a totally connected Gibbs distribution of some spin system on $G$. If $\mu$ is both $u$-marginally bounded and $\eta$-spectrally independent and $n\geq\frac{24d}{u^{2}}(\frac{4\eta}{u^{2}}+1)$, then $\mu$ satisfies the approximate tensorization of entropy with constant
$C_{1}=\frac{18\log(1/u)}{u^{4}}\left(\frac{24d}{u^{2}}\right)^{\frac{4\eta}{u^{2}}+1}.$
Spectral independence and marginal boundedness properties for graphs with constant maximum degree are proven in [CLV23, CLV21] respectively. This shows that the log-Sobolev constant of a graph with bounded degree is $\Omega(n^{-1})$. Having showed that Glauber dynamics chain satisfies desired Log-Sobolev constant and $\|H\|_{P}$, we present the final run time.
Theorem 5.12.
Let $P$ be a Glauber dynamics chain on a graph $G(\mathcal{N},\mathcal{E})$ with a bounded maximum degree $d$ at fugacity parameter $\lambda<\lambda_{c}$ and stationary distribution $\pi$. Then, there exists a short-path algorithm that finds the maximum independent set in $G$ with running time
$\mathcal{O}^{\ast}\left([\pi(E^{\ast})^{-1}]^{\frac{1}{2}-c}\right),$
where $c>0$ is a constant.
Proof.
It is evident that for sparse graphs ($d=\mathcal{O}(1)$), the size of the maximum independent set is $\Theta(n)$. Furthermore, $\log(1/\pi_{\lambda}(E^{\ast}))=\Theta(n)$ due to the hardness of finding the maximum independent sets. As log-Sobolev constant $\omega^{-1}=\mathcal{O}(n)$ for the Glauber dynamics on bounded degree graphs below the critical fugacity $\lambda_{c}$. Therefore, the conditions in Theorem 5.7 holds. Hence, we obtain a super-quadratic speedup over sampling from $\pi_{\lambda}$. ∎
Finally, we show that on a random regular graph, Markov Chain search using Glauber dynamics is better than brute force search.
Proposition 5.13.
Let $\pi_{\lambda}$ be the stationary distribution of Glauber dynamics on a random $d$-regular graph $G\sim\mathcal{G}(n,d)$ with $d\geq d_{0}$. Choose $x^{\ast}$ to be a particular maximum independent set in $G$. Then, for $\lambda=\lambda_{c}$ we have
$\pi(\mathcal{I}^{\ast})\geq 2^{-\kappa n},$
with $\kappa=\frac{2\log(\lambda)\log d_{0}}{d_{0}}-\frac{1}{2d_{0}}-\frac{\log(1+\lambda)}{2}$.
Proof.
The probability of $x^{\ast}$ in $\pi$ is given by
$\pi_{\lambda}(x^{\ast})=\frac{\lambda_{c}^{|x^{\ast}|}}{\sum_{x\in\mathcal{X}}\lambda^{|x|}}.$
We first bound the denominator. To do that, we invoke [Zha09, Theorem 2], which asserts that
$Z(\lambda)=\sum_{x\in\mathcal{X}}\lambda^{|x|}\leq(2(1+\lambda)^{d}-1)^{\frac{n}{2d}},$
for any $d$-regular graph.
For the numerator, we need to bound $|x^{\ast}|$. A known upper bound for the size of the maximum independent set is given by $2\log(d)n/d$ [Bol81]. Combining these, we have
$\pi(x^{\ast})\geq\frac{\lambda^{(2\frac{n}{d}\log d)}}{(2(1+\lambda)^{d}-1)^{\frac{n}{2d}}}\geq\frac{\lambda^{(2\frac{n}{d}\log d)}}{(2(1+\lambda)^{d})^{\frac{n}{2d}}}=2^{n(\frac{2\log(\lambda)\log d}{d}-\frac{1}{2d}-\frac{\log(1+\lambda)}{2})}.$
Now define $f(d)=\frac{2\log(\lambda)\log d}{d}-\frac{1}{2d}-\frac{\log(1+\lambda)}{2}$, where $\lambda=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$. Then, for $d>d_{0}$, we have $f(d)>-1$. ∎
Remark 5.14.
The parameter $\kappa$ from Proposition 5.13 is a decreasing function of $d_{0}$. Therefore, the performance of this algorithm improves as the degree of the graph increases.
5.2.2 Ising Model
Consider the 2-spin Ising model on a graph $G(\mathcal{N},\mathcal{E})$ defined via the Hamiltonian
$H(x)=\sum_{(i,j)\in\mathcal{E}}x_{i}x_{j}+\sum_{j}h_{j}x_{j},$
where the entries of $J$ are interaction coefficients and $h$ defines an external field.
Assume $J_{ij}=1$ for all $(i,j)\in\mathcal{E}$ and $h=\mathbf{0}$. This model corresponds to Gibbs sampling with weights $\pi(x)\propto\exp(-\beta H(x))$, and a Gibbs sample can be prepared by using Glauber dynamics similar to the hardcore model. The only change is the transition probabilities, which can be computed by the marginal distribution $\pi(x_{i}^{t+1}|x^{t}_{\mathcal{N}\setminus\{i\}})$. For simplicity we consider the anti-ferromagnetic model, where having two neighboring sites have the same spin results in lower probability than having the same spin.
Assuming the underlying graph is sparse ($d=\mathcal{O}(1)$), then Glauber dynamics mixes in $\operatorname{poly}(n)$ time for $\beta<\frac{d-2}{d}=\beta_{c}$ [CLV21] at which the uniqueness/non-uniqueness phase transition occurs. We consider the following short path Hamiltonian
$H_{b}(x)=-D(P_{\beta}(x))+bg_{\eta}\left(\frac{H}{|E^{\ast}|}\right),$
where $P_{\beta}$ is the Glauber dynamics transition matrix for $\beta<\beta_{c}$. Similar to the setting of the MIS problem, a block-encoding of $D(P_{\beta})$ can be prepared efficiently.
Proposition 5.15.
The optimum energy of Ising Model Hamiltonian $H$ on a random regular graph $G(\mathcal{N},\mathcal{E})$ satisfies $|E^{\ast}|=\Theta(n)$.
Proof.
[…]
$|\textrm{BW}|=\frac{s+E_{\textrm{gs}}}{2}.$
Using this equality, $E_{\textrm{gs}}=s-2|BW|$. Next, we consider random regular graphs. For sparse random regular graphs $s=\Theta(n)$ and $|\textrm{BW}|=\Theta(n)$ (See [DSW07, COLMS22]). ∎
Theorem 5.16.
Let $P$ be a Glauber dynamics chain on a random regular graph $G(\mathcal{N},\mathcal{E})$ with constant degree $d$ at inverse temperature parameter $\beta<\beta_{c}$ and stationary distribution $\pi_{\beta}$. Then there exists a short-path algorithm that finds the optimum of Ising model Hamiltonian on a random regular graph with running time
$\mathcal{O}^{\ast}\left([\pi_{\beta}(E^{\ast})^{-1}]^{\frac{1}{2}-c}\right),$
where $c>0$ is a constant.
Proof.
As spectral independence and marginal boundedness for Ising model on sparse graphs are proven in [CLV21]. Therefore, Glauber dynamics for Ising model on a regular graph has $\omega^{-1}=\mathcal{O}(n)$. Furthermore, $\log(1/\pi_{\beta}(E^{\ast}))=\Theta(n)$ due to hardness of the problem. Similar to MIS problem, for all $x,x^{\prime}\in\{-1,1\}^{n}$ such that $P(x,x^{\prime})>0$, $\lvert H(x)-H(x^{\prime})\rvert\leq 2d=\mathcal{O}(1)$. By Proposition 5.15, $|E^{\ast}|=\mathcal{O}(n)$. Therefore, the conditions in Theorem 5.7 are satisfied. Hence, we have the super-quadratic sampling over sampling from $\pi_{\beta}$. ∎
5.2.3 Sherrington-Kirkpatrick Model
Consider the (possibly diluted) Sherrington-Kirkpatrick Hamiltonian on a graph $G(\mathcal{N},\mathcal{E})$,
$H(x)=\frac{1}{\sqrt{n}}\sum_{(i,j)\in\mathcal{E}}^{n}g_{ij}x_{i}x_{j},$ (43)
where the interaction coefficients $g_{ij}$ are i.i.d. standard Gaussian random variables. We can Gibbs sample from the following distribution,
$\pi(x)\propto\exp(-\beta H(x))$
using Glauber dynamics.
Lemma 5.17.
Let $P$ be a Glauber dynamics chain for the SK model on a graph $G(\mathcal{N},\mathcal{E})$. Then,
$\Delta_{P}=\mathcal{O}(1).$
Proof.
We first consider hypercube walk. If we flip a spin at random, the sign of each term in $H$ will flip with probability $2/n$. Therefore the energy of each term increases at most by a factor of $1-\frac{4}{n}$ in expectation. Since the ground state energy $|E^{\ast}|=\Theta(n)$, the energy increases at most by constant in expectation. From the definition of Glauber dynamics a bit flip is proposed uniformly, and accepted with probability larger than $\frac{1}{2}$ if $\beta(H(x^{\prime})-H(x))<0$. Therefore, if we are running the Glauber dynamics at some positive finite temperature, moves that increase energy are made with strictly lower probability than the hypercube walk. Thus an uppper bound on stability with respect to the hypercube walk is also a valid upper bound for Glauber dynamics at positive $\beta$. ∎
Lemma 5.18.
Let $P$ be a Glauber dynamics chain for SK model on a graph $G(\mathcal{N},\mathcal{E})$ with constant maximum degree, at inverse temperature $\beta<\beta_{c}$. Then, the log-Sobolev constant satisfies $\omega\geq\Omega(1/(n\log n))$.
Proof.
By [AJK+22, Theorem 12, part (a)], the modified Log Sobolev constant for Glauber dynamics is $\Omega(1/n)$ when $\beta<\beta_{c}$ (See the discussion in Page 11). On a graph with constant bounded degree, the transition probability of Glauber dynamics $\Theta(n^{-1})$. Hence, by Theorem 4.18, the log-Sobolev constant scales as $\Omega(1/(n\log(n)))$. ∎
Theorem 5.19.
Let $P$ be a Glauber dynamics chain for the Sherrington-Kirkpatrick model on a regular graph $G(\mathcal{N},\mathcal{E})$ with a bounded maximum degree $d$ at inverse temperature parameter $\beta<\beta_{c}$ and stationary distribution $\pi_{\beta}$. Then there exists a short-path algorithm that finds the optimal solution of the Sherrington Kirkpatrick Hamiltonian with running time
$\mathcal{O}\left(\operatorname{poly}(n)[\pi_{\beta}(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{c}{\log(n)}\right)}\right),$
where $c>0$ is a constant.
Proof.
We first show that the tail bound holds for SK model. By using proposition 4 in [DPCB23], we know that the number of low energy states with energy smaller than $E^{\ast}(1-\eta)$ is smaller than $2^{\gamma n}$ where $\gamma$ is a constant. By assuming that $\log(1/\pi(E^{\ast}))=\Theta(n)$, we can conclude that the generalized tail bound holds as well. Note that if this assumption fails, then it means that there exists a sub-exponential solver for SK model. Finally, since $\lvert E^{\ast}\rvert=\Theta(n)$ and $\log(1/\pi(E^{\ast}))=\Theta(n)$ for the SK Model by using Lemma 5.18 and Theorem 4.9, we have $\gamma=\mathcal{O}(1)$. Thus, by Theorem 4.15, we have $b=\mathcal{O}(1/\log(n))$. Since $\Delta$ is constant by Lemma 5.17, the total runtime scales as $(\pi_{\beta}(E^{\ast}))^{-1\left(\frac{1}{2}-\frac{c}{\log(n)}\right)}$ due to Theorem 3.3. ∎
We conclude this section by demonstrating that Markov Chain search using Glauber dynamics at a positive inverse-temperature is faster than unstructured search.
Lemma 5.20.
Let $\pi$ be the Gibbs distribution corresponding to a cost function $H\colon\{0,1\}^{n}\mapsto\mathbb{R}$ at some positive inverse temperature $\beta>0$. We also assume that the cost function is concentrated away from its optimum, i.e., the number of states with cost greater than $(1-\eta)E^{\ast}$ is lower than $2^{-\gamma n}$.
Proof.
We first observe due to the assumption on concentration that the partition function $Z(\beta)\leq 2^{n}[(1-2^{-\gamma n})\exp(-\beta(1-\eta)E^{\ast})+2^{-\gamma n}\exp(-\beta E^{\ast})]$. As a consequence,
$\frac{\pi(E^{\ast})}{2^{-n}}$ $\geq\frac{\exp(-\beta E^{\ast})}{(1-2^{-\gamma n})\exp(-\beta(1-\eta)E^{\ast})+2^{-\gamma n}\exp(-\beta E^{\ast})}$ (44)
$\geq\frac{1}{0.5\exp(-\beta\eta|E^{\ast}|)+2^{-\gamma n}}$ (45)
$=\Omega\left(\max\left(2^{\gamma n},\exp(\beta\eta|E^{\ast}|)\right)\right).$ (46)
6 Numerical Results
We perform numerical evaluations to empirically verify our findings. We focus on the constrained problems studied in this work, including (MaxBisection), MaxCut with a Hamming weight constraint $k=o(n)$ (MaxCut-Hamming), and MIS with a penalized objective. For (MaxBisection), we take $n$ to be even since $k=\frac{n}{2}$. For (MaxCut-Hamming), we take $k=\lfloor\sqrt{n}\rfloor$. For MIS, we take the penalty factor to be $n$, such that no energy reduction from constraint violation can justify the penalty. For all three problems, we generate 100 random unweighted graphs for each $n$ from the Erdős–Rényi model with the probability of each edge existing to be $\frac{2\ln n}{n}$. The constant factor 2 is chosen to ensure a reasonable graph density at the scale we cover. We set $\eta=0.5$ in all experiments.
To improve the scalability of our numerical experiments, we construct $H_{b}$ as a sparse matrix in the compressed sparse row format and employ a GPU-accelerated iterative eigensolver to compute only the two smallest eigenvalues and the corresponding eigenvectors. For (MaxBisection) and (MaxCut-Hamming), which are explicitly constrained, we can scale up further by projecting $H_{b}$ onto the space spanned by all feasible states. The dimension of the computational space then drops from $2^{n}$ to $\binom{n}{k}$. With these efforts, we obtained results with up to 30 qubits for (MaxCut-Hamming).
Figure 1: Empirical selection of $b$. A Quartiles of $b$ values that minimize the effective runtime of the algorithm for (MaxCut-Hamming). As $n$ increases, the runtime-optimal $b$ converges to a range approximately between 0.8 and 1.2. The red dot line shows the converged value of $b\approx 0.78$ of phase transition where the overlap with the initial state crosses $0.99$. B Quartiles of $b$ values that minimize the spectral gap for (MaxCut-Hamming). For most instances tested, the spectral gap is minimized when $b$ is larger than the phase transition value, rendering the phase transition $b$ a safe choice. C The overlap values with the initial state and the ground state (optimal solution) for one $n=30$ (MaxCut-Hamming) instance with varying $b$. The dotted verticle line denotes the phase transition $b$.
Figure 2: The inverse of the ground state overlap versus the feasible space size $\binom{n}{k}$ for (MaxCut-Hamming) with $n$ varying from 10 to 30 and $b=$ 0.78. The worst-case instances are fitted using an exponential function with base $\binom{n}{k}$ with an error bar denoting one standard deviation of the fitted exponent. The 95% confidence interval on the fitted exponent is $[0.391,0.408]$.
First, we want to identify what values of $b$ are practically appropriate. In [DPCB23], the authors numerically show that the original short-path algorithm works well for the 3-spin problem for $b$ up to around 0.8, which is much larger than the theoretical bound of $b\leq 1.02\times 10^{-4}$. Here, we show that a similar observation can be made for the constrained and penalized cases. For (MaxCut-Hamming), Figure 1 A shows the quartiles of $b$ values that minimize the effective runtime (Equation 9) of the algorithm. Note that the $b$ values are hard-capped at 1.25 and may be higher. We see that as $n$ increases, the optimal $b$ converges to a range approximately between 0.8 and 1.2. However, when choosing the value of $b$ that needs to work for all instances, we want a conservative value that avoids encountering the possibly superexponentially small spectral gap. For this purpose, we identify the value of $b$ at which the phase transition occurs. We numerically characterize the phase transition point by the overlap of |$\ket{\psi_{b}}$ with the initial state dropping below 0.99. An example of the overlap with varying $b$ is shown in Figure 1 C. Empirically, we observe that the phase transition $b$ converges to around $0.78$ as $n$ increases. In Figure 1 B, we plot the quartiles of $b$ values that minimize the spectral gap. For most instances, the spectral gap is minimized when $b$ is greater than the phase transition value $\approx 0.78$. Therefore, we expect a ubiquitous value of phase transition to work for a (MaxCut-Hamming) instance with high probability.
We then fit the worst-case runtime to empirically demonstrate the super-Grover speedup. Although the inverse of the spectral gap term in the runtime (Equation 9) has a $\mathcal{O}(\operatorname{poly}(n))$ complexity, it may still affect the exponential fitting at the scale of numerical experiments. Thus, we use the inverse of ground state overlap $\lvert\langle\psi_{b}|z\rangle\rvert^{-1}$, the only exponential growth term in the runtime, to fit the asymptotic speedup. In Figure 2, we set $b$ to be 0.78 and plot the inverse of ground state overlap $\lvert\langle\psi_{b}|z\rangle\rvert^{-1}$ of all (MaxCut-Hamming) instances with respect to $\binom{n}{k}$, the size of the feasible space. We fit the worst-case instances using an exponential function with base $\binom{n}{k}$, the exponent of which is equivalent to the factor $a$ in $2^{an}$ for the unconstrained case. The error bar of the fitted line denotes one standard deviation of the fitted exponent. We see the empirical $b$ values give a super-Grover speedup, which is much better than the theoretically guaranteed bounds.
In Figure 3, we show the empirical worst-case scaling for all three problems with different choices of $b$. A proper selection of $b$ yields a super-Grover speedup across all examined problems. Conversely, when $b$ is excessively high, the algorithm may encounter a small spectral gap in the worst case. To demonstrate this, we use the runtime (Equation 9, which includes the inverse of the spectral gap term) as the metric and observe that the quality of the fitting degrades. Our numerics lead to two interesting conceptual observations: firstly, as observed also by [DPCB23] the optimal choices of $b$ are well beyond what is predicted by the theoretical analysis. Secondly, in the case of (MaxCut-Hamming) we numerically observe an advantage over quadratic speedup that does not decay with $n$ which is beyond the current theoretical analysis and indicates that the runtime of the long jump can possibly be characterized through weaker conditions than $\Delta_{p}$ smoothness. Finally, we confirm in our setting that is indeed reasonable to make the choice of $b$ by choosing the largest such value that allows for large overlap with the ground state. If the value of this critical $b$ asymptotes quickly as a function of $n$ this suggests a numerical mechanism for the development of efficient short-path algorithms.
Figure 3: Empirical fitting of the inverse overlap of the ground state $\lvert\langle\psi_{b}|z\rangle\rvert^{-1}$ and the runtime 9 of (MaxCut-Hamming), (MaxBisection), and MIS with difference choices of $b$. For the left column (MaxCut-Hamming), we use data with $n$ ranging from 10 to 30. Top: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.383,0.402]$, indicating there could exist $b$ that is better than the phase transition one in Figure 2; Bottom: $b=1$ the fitted exponent is 0.348 with 95% confidence interval $[0.271,0.426]$. For the middle column (MaxBisection), we use data with $n$ ranging from 16 to 22. Top: $b=0.7$ the fitted exponent is 0.444 with 95% confidence interval $[0.436,0.452]$; Bottom: $b=1$ the fitted exponent is 0.873 with 95% confidence interval $[0.842,0.905]$. For the right column (MIS), we use data with $n$ ranging from 10 to 21. Top: $b=0.6$ the fitted exponent is 0.400 with 95% confidence interval $[0.386,0.415]$; Bottom: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.122,0.663]$.
Acknowledgements
The authors thank Aram Harrow, Shree Hari Sureshbabu, and Jiayu Shen for insightful discussions and feedback, and their colleagues at the Global Technology Applied Research center of JPMorganChase for their support and helpful discussions.
Appendix A Technical Details for Generalized Short Path Framework
Lemma A.1 ($\theta$ short-path $\implies$ $\theta$ Spectral Gap Bound, adapted from Proposition 5 of [DPCB23]).
If $H_{b}$ satisfies the $\theta$ short-path condition, then the spectral gap of $H_{b}$ is at least $\theta$, i.e., all excited states have energy at least $-1+\theta$.
Proof.
Recall by construction, the ground state energy of $H_{b}$ is at most $-1$. Note that $\theta\leq 1$. Suppose, in order to arrive at a contradiction, that there are at least two orthogonal eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, with energy strictly below $-1+\theta$. Since the short-path condition is satisfied, at least one of $|\psi_{1}\rangle$ or $|\psi_{2}\rangle$ has nonzero overlap with $|\pi\rangle$. Without loss of generality, assume this is the case for $|\psi_{2}\rangle$.
Now consider the state
$|\psi^{\prime}\rangle=\left(1+\frac{\lvert\langle\sqrt{\pi}|\psi_{1}\rangle\rvert^{2}}{\lvert\langle\pi|\psi_{2}\rangle\rvert^{2}}\right)^{-1/2}\left(|\psi_{1}\rangle-\frac{\langle\sqrt{\pi}|\psi_{1}\rangle}{\langle\sqrt{\pi}|\psi_{2}\rangle}|\psi_{2}\rangle\right),$
which is orthogonal to $|\sqrt{\pi}\rangle$. Since $\textup{{GSE}}\left(\Pi_{\perp}H_{b}\Pi_{\perp}\right)\leq\langle\psi^{\prime}|H_{b}|\psi^{\prime}\rangle$, and
$\langle\psi^{\prime}|\Pi_{\perp}H_{b}\Pi_{\perp}|\psi^{\prime}\rangle$ $=\langle\psi^{\prime}|H_{b}|\psi^{\prime}\rangle$
$<\left(1+\frac{\lvert\langle\sqrt{\pi}|\psi_{1}\rangle\rvert^{2}}{\lvert\langle\sqrt{\pi}|\psi_{2}\rangle\rvert^{2}}\right)^{-1}\left[(-1+\theta)+\frac{\lvert\langle\sqrt{\pi}|\psi_{1}\rangle\rvert^{2}}{\lvert\langle\sqrt{\pi}|\psi_{2}\rangle\rvert^{2}}(-1+\theta)\right]\leq-1+\theta,$
we get a contradiction. ∎
Lemma A.2 (Runtime bound by approximate projector).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain with $|\mathcal{X}|=V$. Define
$\mathcal{P}_{\ell}:=\left(\frac{H_{b}}{\lvert E_{b}\rvert}\right)^{\ell},$
and let $\omega$ be the log-Sobolev constant of $D(P)$. For either all even or all odd $\ell$ we have
$\left\langle\sqrt{\pi}\middle|\mathcal{P}_{\ell}\middle|z\right\rangle-V\left(1-\frac{\omega}{2}\right)^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z\right\rangle.$
Proof.
Note that $E_{b}\leq-1$. By Lemma 4.12 all excited states have energy at least $-1+\frac{\omega}{2}$.
Since $\omega$ is the log-Sobolev constant of $D(P)$ (lower bounding its spectral gap $\delta$ as defined in (5)) and $g_{\eta}(H)$ is negative definite, at most two eigenvectors of $H_{b}$ have have magnitude $>1-\frac{\omega}{2}$, the ground state and highest-energy state. Thus,
$\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z\rangle$ $=\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+\sum_{|E_{b}^{\prime}|>1-\frac{\omega}{2}}\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle+\sum_{|E_{b}^{\prime}|\leq 1-\frac{\omega}{2}}\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle$
$\leq\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle+(V-2)(1-\frac{\omega}{2})^{\ell}.$
Supposing $\ell$ takes a value such that the middle term is non-positive, we obtain
$\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z\rangle-V(1-\frac{\omega}{2})^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle.$
Lemma A.3 (ground-state energy shift bound, adapted from Proposition 6 of [DPCB23]).
Suppose that $\gamma$-spectral density and $\frac{\omega}{2}$-short-path condition hold, then
$\lvert E_{b}\rvert<1+\frac{4(\pi(E^{\ast}))^{\gamma}}{\omega}.$ (47)
Proof.
In the proof of Proposition 6 of [DPCB23], the authors showed that
$E_{b}$ $=\langle\psi_{b}|H_{b}|\psi_{b}\rangle=-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-b^{2}\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle,$ (48)
where
$W_{b}$ $:=(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp})^{-1}=\Pi_{\perp}(H_{b}-E_{b})^{-1}\Pi_{\perp}$
$\Pi_{\perp}$ $:=I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|.$
Recall the short-path condition
$\textup{{GSE}}(\Pi_{\perp}H_{b}\Pi_{\perp})\geq-1+\frac{\omega}{2}.$
Thus, with $E_{b}\leq-1$,
$\textup{{GSE}}(\Pi H_{b}\Pi)\geq-1+\frac{\omega}{2}\geq E_{b}+\frac{\omega}{2}$
$\implies\textup{{GSE}}(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp}))\geq\omega/2,$
so $\lVert W_{b}\rVert_{2}\leq\frac{2}{\omega}$ and as a consequence,
$\langle\sqrt{\pi}|G_{\eta}(\Pi(H_{b}-E_{b})\Pi)^{-1}G_{\eta}|\sqrt{\pi}\rangle\leq\frac{2}{\omega}\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle.$
From here, applying Equation (48) yields
$E_{b}\geq-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-\frac{2b^{2}}{\omega}\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle.$
By spectral density and definition of $G_{\eta}$:
$-\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle$ $\leq(\pi^{\ast})^{\gamma}$
$\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle$ $\leq(\pi^{\ast})^{\gamma}.$
Using $0\leq b\leq 1$, $\pi^{\ast}<1,\omega<1$, we have
$E_{b}\geq-1-(\pi^{\ast})^{\gamma}\left(b+b^{2}\frac{2}{\omega}\right)>-1-\frac{4(\pi^{\ast})^{\gamma}}{\omega}.$
Next we provide an alternative bound on the energy shift. At a high level, the goal is to show that the small ground state energy shift condition is satisfied for any point we can efficiently go to in the short jump. It removes the need for the short-path condition. Note, heuristically, under this large ground state overlap condition, the short-path condition reduces to the large gap condition:
$\textup{{GSE}}((I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|)H_{b}(I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|))-E_{b}\approx\textup{{GSE}}((I-|\psi_{b}\rangle\langle\psi_{b}|)H_{b}(I-|\psi_{b}\rangle\langle\psi_{b}|))-E_{b}=\text{Gap}(H_{b}).$
Lemma A.4 ($\Delta_{P}(\eta)$ stable $\iff\alpha_{P}$ subdepolarizing ).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain. The pair $(\mathcal{M},H)$ satisfies the $\Delta_{P}(\eta)$-stable if and only if it satisfies $\alpha_{P}$-subdepolarizing. They are related by the equation $\alpha_{P}=\frac{\Delta_{P}(\eta)}{\lvert E^{\ast}\rvert(1-\eta)}$.
Proof.
The $\implies$ direction follows mostly from the proof of Proposition 3 in [DPCB23] By definition of $g_{\eta}$, $f$ is monotonically non-decreasing. From [DPCB23, Proposition 12], $\prod_{t=1}^{T}f(c_{t}x)$ is also a convex function. Thus
$\sum_{y}P(x,y)\prod_{t=1}^{T}f\left(\frac{c_{t}H(y)}{E^{\ast}}\right)$ $\geq\prod_{t=1}^{T}f\left(\sum_{y}P(x,y)\frac{c_{t}H(y)}{E^{\ast}}\right)$
$\geq\prod_{t=1}^{T}f\left(\frac{c_{t}H(x)}{E^{\ast}}\left(1+\frac{\Delta}{H(x)}\right)\right).$
Recall that $g_{\eta}(z)$ is zero for all $z\geq(1-\eta)E^{*}$. If for some $t$, $\frac{c_{t}H(x)}{E^{*}}>(1-\eta)E^{*}$, then the whole product is zero, as in this case
$\frac{c_{t}H(x)}{E^{\ast}}\left(1+\frac{\Delta}{H(x)}\right)>(1-\eta)E^{*}.$
Thus, the stated hypothesis is satisfied trivially. If
$\frac{H(x)}{E^{*}}<\frac{(1-\eta)E^{*}}{c_{t}}\leq(1-\eta)E^{*}$
for all $t$, then
$\frac{1}{H(x)}>-\frac{1}{(1-\eta)\lvert E^{*}\rvert}.$
By monotonicity of $f$
$\prod_{t=1}^{T}f\left(\frac{c_{t}H(x)}{E^{\ast}}\left(1+\frac{\Delta}{H(x)}\right)\right)\geq\prod_{t=1}^{T}f\left(\frac{c_{t}H(x)}{E^{\ast}}\left(1-\frac{\Delta}{(1-\eta)\lvert E^{*}\rvert}\right)\right).$
For the $\Leftarrow$ direction, consider taking $c_{t}\rightarrow 0,t\neq 0$ and $c_{1}\rightarrow 1$. By continuity we have
$\sum_{y}P(x,y)f\left(\frac{H(y)}{E^{\ast}}\right)\geq f\left(\frac{(1-\alpha_{P})H(x)}{E^{\ast}}\right)$
We can suppose $H(x)<-(1-\eta)\lvert E^{*}\rvert$, since otherwise the right-hand side is zero, and so our choice of $\Delta$ in terms of $\alpha_{P}$ clearly works. Thus,
$\sum_{y}P(x,y)f\left(\frac{H(y)}{E^{\ast}}\right)\geq f\left(\frac{H(x)+\alpha_{P}(1-\eta)\lvert E^{*}\rvert}{E^{*}}\right).$
so by definition of $f$
$\sum_{y}P(x,y)g_{\eta}\left(\lvert E^{*}\rvert^{-1}H(y)\right)\leq g_{\eta}\left(\lvert E^{*}\rvert^{-1}(H(x)+\alpha_{P}(1-\eta)\lvert E^{*}\rvert)\right).$
Lemma A.5 (Upper bounds on $\Delta(\eta)$).
Let $\eta\in[0,1)$. If $H$ has $P$ pseudo-Lipschitz norm $\lVert H\rVert_{P}$, then
$\sqrt{\lVert H\rVert_{P}}\geq\Delta_{P}(\eta).$
Furthermore, if
[…]
then for $\eta\in[0,1)$
$\tilde{\Delta}_{P}\geq\Delta_{P}(\eta).$
Proof.
Note that if have the stronger condition, for some $\tilde{\Delta}_{P}>0$,
[…]
which may actually be easier to show, then we also have $\Delta_{P}(\eta)$ stability with $\Delta_{P}(\eta)\leq\tilde{\Delta}_{P}$, since by concavity of $h_{\eta}=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$ and Jensen’s inequality:
[…]
Also from the above, it is a simple consequence of Jensen’s inequality that we can take $\Delta_{P}(\eta)$ to be the $\sqrt{\lVert H\rVert_{P}}$:
$\sqrt{\lVert\psi\rVert_{P}}$ $=\sqrt{\max_{x\in\mathcal{X}}\sum_{y}P(y,x)(H(x)-H(y))^{2}}$
$\geq\max_{x\in\mathcal{X}}\sum_{y}P(y,x)\lvert H(x)-H(y)\rvert$
$\geq\tilde{\Delta}_{P}$
$\geq\Delta_{P}(\eta),\quad\forall\eta\in[0,1).$
Appendix B Technical Details for MaxCut Hamming and MaxBisection
Lemma B.1.
Let $G(\mathcal{N},\mathcal{E})$ be drawn from the Erdős-Rényi ensemble $\mathcal{G}\left(n,\frac{p}{n-1}\right)$ for a constant $p$. Consider the objective of (MaxCut-Hamming):
$H(x):=-\frac{1}{2}\sum_{i<j}e_{ij}(1-x_{i}x_{j}).$
Then, with probability at least $1-\delta$ over the graph, it follows:
$\left\lvert-\mathbb{E}_{\pi}H(x)-\frac{pk(n-k)}{n}\right\rvert\leq C\log(\delta^{-1})\frac{2\sqrt{p}k(n-k)}{n\sqrt{2(n-1)}},$
where $C>0$ is an arbitrary constant, and $\pi$ is the uniform distribution of Hamming-weight $k$ strings.
Proof.
The shift required to ensure that the mean over in-constraint strings is zero is given by
$-\mathbb{E}_{\pi}H(x)=\sum_{i<j}e_{ij}\mathbb{E}_{\pi}\frac{1}{2}(1-x_{i}x_{j})=\sum_{i<j}e_{ij}\Pr[x_{i}\neq x_{j}].$
[…]
$\Pr[x_{i}\neq x_{j}]=\frac{2{n-2\choose k-1}}{{n\choose k}}=\frac{2k(n-k)}{n(n-1)}.$
Suppose edge creation probability is $\frac{p}{n}$. Since $\sum_{i<j}e_{ij}$ a Binomial random variable $B\left(\binom{n}{2},\frac{p}{n}\right)$, applying the Chernoff bound asserts that with probability at least $1-\delta$, we have:
$\left\lvert\sum_{i<j}e_{ij}-\binom{n}{2}\frac{p}{n}\right\rvert\leq C\log(\delta^{-1})\sqrt{\binom{n}{2}\frac{p}{n}\left(1-\frac{p}{n}\right)},$
where $C>0$ is an arbitrary constant.
Accordingly, with probability $1-\delta$, it follows:
$\left\lvert\mathbb{E}_{\pi}H(x)-\binom{n}{2}\frac{2k(n-k)}{n(n-1)}\frac{p}{n}\right\rvert$ $\leq C\log(\delta^{-1})\frac{2k(n-k)}{n(n-1)}\sqrt{\binom{n}{2}\frac{p}{n}\left(1-\frac{p}{n}\right)}$
$\implies\left\lvert\mathbb{E}_{\pi}H(x)-\frac{pk(n-k)}{n}\right\rvert$ $\leq C\log(\delta^{-1})\frac{2\sqrt{p}k(n-k)}{n\sqrt{2(n-1)}}.$
Thus, for $k=\frac{n}{2}$ one has
$\left\lvert\mathbb{E}_{\pi}H(x)-\frac{pn}{4}\right\rvert\leq C\log(\delta^{-1})\frac{\sqrt{p}n}{2\sqrt{2(n-1)}}\implies\left\lvert\frac{\mathbb{E}_{\pi}H(x)}{n}-\frac{p}{4}\right\rvert\leq C\log(\delta^{-1})\frac{\sqrt{p}}{2\sqrt{2(n-1)}},$
from which one can conclude
$-\frac{\mathbb{E}_{\pi}H(x)}{n}=\frac{p}{4}(1+o(1))$ (49)
with high probability. Thus for $k=n/2$, $\lvert E^{\ast}\rvert=\Theta(n)$ so the cost function only gets shifted by constant to make it mean zero. More generally, for $k=o(n)$, with high probability
$-\mathbb{E}_{\pi}H(x)=pk(1+o(1)).$
Lemma B.2.
Let $H$ be the cost function of (MaxCut-Hamming):
$H(x):=-\sum_{i<j}e_{ij}\frac{(1-x_{i}x_{j})}{2},$
and $P$ be the transition matrix for the transposition walk on the space Hamming-weight $k$ bitstrings. Then, for an arbitrary graph $G$ with $\lvert\mathcal{E}\rvert$ edges and Hamming-weight $k$ MaxCut $\mathcal{C}^{*}_{k}$, it follows:
$\frac{\mathcal{C}^{*}_{k}(n-2)}{k(n-k)}\geq\mathbb{E}_{y\sim x}[H(y)]-H(x)\geq\frac{\mathcal{C}^{*}_{k}(n-2)}{k(n-k)}-\left(\lvert\mathcal{E}\rvert-\mathcal{C}^{*}_{k}\right)\max\left\{\frac{2}{k},\frac{2}{n-k}\right\},$
for every $x\in\{u\in\{-1,1\}^{n}:|u|=k\}$.
Proof.
Fix a graph $G$, and consider an arbitrary $x\in\{-1,1\}^{n}$ such that $\lvert x\rvert=k$ corresponding to number of $+1$’s. Swaps only occur between $x_{i}$ and $x_{j}$ that are different. All one step transitions denoted by $\sim$ are implied to be under the transposition walk.
$H(x)$ $=-\sum_{i<j}e_{ij}\frac{(1-x_{i}x_{j})}{2}=-\sum_{i<j~{}|~{}x_{i}\neq x_{j}}e_{ij},$
$\mathbb{E}_{y\sim x}[H(y)]$ $=-\sum_{i<j}e_{ij}\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\right],$
$\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\right]$ $=\mathbb{P}[y_{i}\neq y_{j}|x].$
For a given fixed $x\in\{-1,1\}^{n}$ satisfying $\lvert x\rvert=k$, we know there are $k(n-k)$ pairs of indices such that $x_{i}\neq x_{j}$, $\binom{n-k}{2}$ indices where $x_{i}=x_{j}=-1$, and $\binom{k}{2}$ indices where $x_{i}=x_{j}=1$. Thus we can decompose the expectation as follows:
$\mathbb{E}_{y\sim x}[H]=$ $-\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}\neq x_{j}]-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}=x_{j}=1]$ (50)
$-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}=x_{j}=-1].$
Using the following facts:
$\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}\neq x_{j}]=\frac{(k-1)(n-k-1)+1}{k(n-k)}$
$\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}=x_{j}=1]=\frac{2}{k}$
$\mathbb{P}_{y\sim x}[y_{i}\neq y_{j}|x_{i}=x_{j}=-1]=\frac{2}{n-k}.$
the expression (50) simplifies to
$\mathbb{E}_{y\sim x}[H]=$ $-\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\frac{(k-1)(n-k-1)+1}{k(n-k)}-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\frac{2}{k}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\frac{2}{n-k}.$
As a consequence, for all $x$ with Hamming weight $k$,
$\mathbb{E}_{y\sim x}[H]-H(x)$
$=-\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\left(\frac{(k-1)(n-k-1)+1}{k(n-k)}-1\right)-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\frac{2}{k}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\frac{2}{(n-k)}$
$=\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\frac{n-2}{k(n-k)}-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\frac{2}{k}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\frac{2}{(n-k)}.$
Thus for any $x\in\{-1,1\}^{n}$ and any graph $G$:
$\mathbb{E}_{y\sim x}[H]\leq H(x)+\frac{-H(x)(n-2)}{k(n-k)},$
We have:
$\forall x,\mathbb{E}_{y\sim x}[H]\leq H(x)+\frac{\mathcal{C}^{*}_{k}(n-2)}{k(n-k)}.$
For the lower bound, for every $x\in\{u\in\{-1,1\}^{n}:|u|=k\}$, we have:
$\mathbb{E}_{y\sim x}[H(y)]-H(x)\geq\frac{\mathcal{C}^{*}_{k}(n-2)}{k(n-k)}-\left(\lvert\mathcal{E}\rvert-\mathcal{C}^{*}_{k}\right)\max\left\{\frac{2}{k},\frac{2}{n-k}\right\},$
where $\lvert\mathcal{E}\rvert$ is the number of edges. ∎
Lemma B.3.
For the MaxCut-Hamming Hamiltonian $H$, the pseudo Lipschitz constant $\lVert H\rVert_{P}$ under the transposiiton walk is $\mathcal{O}(1)$ with high probability.
Proof.
Recall
[…]
so we can consider
[…]
The only new term is $\mathbb{E}_{y\sim x}[H(y)^{2}]$, which requires us to look at terms like:
[…]
$y$ is $x$ after a single random transposition. We can put these terms in groups based on:
  1. 1.
    $x_{j}=x_{i}=x_{r}=x_{s}=\pm 1$, the term is always zero
  2. 2.
    $x_{j}=x_{i}=x_{r}\neq x_{s}=\pm 1$
    $-1:\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{2(n-k-1)}{k(n-k)}$
    $+1:\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{2(k-1)}{k(n-k)}$
  3. 3.
    $x_{j}=x_{i}\neq x_{r}=x_{s}=\pm 1$ the term is zero unless an element of $\{i,j\}$ is swapped with an element of $\{r,s\}$, giving
    ±$\pm 1:\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{4}{(n-k)k}$
  4. 4.
    $x_{j}\neq x_{i}=x_{r}\neq x_{s}=\pm 1$,
    $\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{(n-k-2)(k-2)+2}{k(n-k)}=\frac{k(n-k)-2(n-3)}{k(n-k)}$
Let $\hat{y}_{ij}=\frac{1-y_{i}y_{j}}{2}$, then expanding:
$H(y)$ $=-\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\hat{y}_{ij}-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\hat{y}_{ij}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\hat{y}_{ij}$
$\mathbb{E}_{y\sim x}[H(y)^{2}]$ $=\mathbb{E}_{y\sim x}\left\{\left(\sum_{i<j~{}:~{}x_{i}\neq x_{j},}e_{ij}\hat{y}_{ij}\right)^{2}\right\}$
$+2\mathbb{E}_{y\sim x}\left\{\left(\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\hat{y}_{ij}\right)\left(\sum_{r<s~{}:~{}x_{r}\neq x_{s}}e_{ij}\hat{y}_{ij}\right)\right\}$
$+2\mathbb{E}_{y\sim x}\left\{\left(\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\hat{y}_{ij}\right)\left(\sum_{r<s~{}:~{}x_{r}\neq x_{s}}e_{ij}\hat{y}_{ij}\right)\right\}$
$+2\mathbb{E}_{y\sim x}\left\{\left(\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\hat{y}_{ij}\right)\left(\sum_{r<s~{}:~{}x_{r}=x_{s}=1}e_{ij}\hat{y}_{ij}\right)\right\},$
where we have eliminated the square of the $x_{i}=x_{j}=\pm 1$ terms since they fall into group $1$. Expanding further and passing the expectations through:
$\mathbb{E}_{y\sim x}[H(y)^{2}]$ $=\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\mathbb{E}_{y\sim x}\hat{y}_{ij}$
$\quad+2\sum_{i<j,r<s,(i,j)\neq(r,s)~{}:~{}x_{i}\neq x_{j},~{}x_{r}\neq x_{s}}e_{ij}e_{rs}\mathbb{E}_{y\sim x}[\hat{y}_{ij}\hat{y}_{rs}]$
$\quad+2\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}\neq x_{s}}e_{ij}e_{rs}\mathbb{E}_{y\sim x}[\hat{y}_{ij}\hat{y}_{rs}]$
$\quad+2\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=-1,x_{r}\neq x_{s}}e_{ij}e_{rs}\mathbb{E}_{y\sim x}[\hat{y}_{ij}\hat{y}_{rs}]$
$\quad+2\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}=x_{s}=-1}e_{ij}e_{rs}\mathbb{E}_{y\sim x}[\hat{y}_{ij}\hat{y}_{rs}].$
Next we compute the expectations by identifying which group they belong to:
$\mathbb{E}_{y\sim x}[H(y)^{2}]$ $=\frac{k(n-k)-(n-2)}{k(n-k)}\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}$
$+\frac{2k(n-k)-4(n-3)}{k(n-k)}\sum_{i<j,r<s,(i,j)\neq(r,s)~{}:~{}x_{i}\neq x_{j},~{}x_{r}\neq x_{s}}e_{ij}e_{rs}$
$+\frac{4(n-k-1)}{k(n-k)}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}\neq x_{s}}e_{ij}e_{rs}$
$+\frac{4(k-1)}{k(n-k)}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=-1,x_{r}\neq x_{s}}e_{ij}e_{rs}$
$+\frac{8}{(n-k)k}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}=x_{s}=-1}e_{ij}e_{rs}.$
The only component to clarify is the first sum, which follows from a calculation in the previous lemma
[…]
For the $2\mathbb{E}_{y\sim x}[H]-H(x)$ part, we use the same computations from the previous lemma, where we computed $\mathbb{E}_{y\sim x}[H]-H(x)$:
$2\mathbb{E}_{y\sim x}[H]-H(x)$
$=-\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\left(\frac{2(k-1)(n-k-1)+2}{k(n-k)}-1\right)-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\frac{4}{k}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\frac{4}{(n-k)}$
$=\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}\frac{-k(n-k)+2(n-2)}{k(n-k)}-\sum_{i<j~{}:~{}x_{i}=x_{j}=1}e_{ij}\frac{4}{k}-\sum_{i<j~{}:~{}x_{i}=x_{j}=-1}e_{ij}\frac{4}{(n-k)},$
where recall that
[…]
Then we compute $-H(x)[2\mathbb{E}_{y\sim x}[H]-H(x)]$:
$-H(x)[2\mathbb{E}_{y\sim x}[H]-H(x)]$
$=\sum_{i<j:~{}x_{i}\neq x_{j}}e_{ij}\frac{-k(n-k)+2(n-2)}{k(n-k)}$
$+\sum_{i<j,r<s,(i,j)\neq(r,s):~{}x_{i}\neq x_{j},x_{r}\neq x_{s}}e_{rs}e_{ij}\frac{-2k(n-k)+4(n-2)}{k(n-k)}$
$-\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}\neq x_{s}}e_{rs}e_{ij}\frac{4}{k}$
$-\sum_{i<j,r<s:~{}x_{i}=x_{j}=-1,x_{r}\neq x_{s}}e_{rs}e_{ij}\frac{4}{(n-k)}.$
We can now put all the expressions together to get:
$\mathbb{E}_{y\sim x}[H(y)^{2}]-H(x)(2\mathbb{E}_{y\sim x}[H(y)]-H(x))$ $=\frac{(n-2)}{k(n-k)}\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}$ (52)
$\quad+\frac{4}{k(n-k)}\sum_{i<j,r<s,(i,j)\neq(r,s)~{}:~{}x_{i}\neq x_{j},~{}x_{r}\neq x_{s}}e_{ij}e_{rs}$ (53)
$\quad-\frac{4}{k(n-k)}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}\neq x_{s}}e_{ij}e_{rs}$
$\quad-\frac{4}{k(n-k)}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=-1,x_{r}\neq x_{s}}e_{ij}e_{rs}$
$\quad+\frac{8}{k(n-k)}\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}=x_{s}=-1}e_{ij}e_{rs}.$ (54)
Let $X$ be a Binomial random variable $\mathcal{B}(M,q)$, then
$\mathbb{P}[X^{2}\geq(1+\delta)^{2}\mathbb{E}[X^{2}]]\leq e^{-\frac{\delta^{2}Mq}{2+\delta}}.$ (55)
This follows simply from Jensen’s inequality
$\mathbb{P}[X^{2}\geq(1+\delta)^{2}\mathbb{E}[X^{2}]]$ $\leq\mathbb{P}[X^{2}\geq(1+\delta)^{2}(\mathbb{E}[X])^{2}]$
$=\mathbb{P}[X\geq(1+\delta)\mathbb{E}[X]]$
$\leq e^{-\frac{\delta^{2}\mathbb{E}[X]}{2+\delta}},$
where the last inequality is the multiplicative Chernoff bound. Also for $X$ and $Y$ being independent Binomials we have that:
$\mathbb{P}[XY\geq(1+\delta)\mathbb{E}[XY]]\leq 2e^{-\frac{\delta^{2}\min(\mathbb{E}[X],\mathbb{E}[Y])}{2+\delta}},$
which follows from
$\mathbb{P}[XY\geq(1+\delta)\mathbb{E}[XY]]\leq\mathbb{P}[X\geq(1+\delta/2)\mathbb{E}[X]\vee Y\geq(1+\delta/2)\mathbb{E}[Y]],$
which follows from
$(X\geq(1+\delta/3)\mathbb{E}[X])\wedge(Y\geq(1+\delta/3)\mathbb{E}[Y])$ $\implies XY\leq(1+\delta)\mathbb{E}[X]\mathbb{E}[Y]$
$=XY\leq(1+\delta)\mathbb{E}[XY],$
so union bound gives the desired result.
Thus, we can assume with high probability all of the sums of Bernoullis in Equation (52) are constant factors from their means. Their means are
  1. 1.
    $\mathbb{E}[\sum_{i<j~{}:~{}x_{i}\neq x_{j}}e_{ij}]\frac{pk(n-k)}{n-1}\asymp pk$
  2. 2.
    $\mathbb{E}[\sum_{i<j,r<s,(i,j)\neq(r,s)~{}:~{}x_{i}\neq x_{j},~{}x_{r}\neq x_{s}}e_{ij}e_{rs}]=[\frac{p}{n-1}k(n-k)]^{2}\asymp p^{2}k^{2}$
  3. 3.
    $\mathbb{E}[\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}\neq x_{s}}e_{ij}e_{rs}]=[\frac{p}{n-1}]^{2}\binom{k}{2}k(n-k)\asymp p^{2}\frac{k^{3}}{n}$
  4. 4.
    $\mathbb{E}[\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=-1,x_{r}\neq x_{s}}e_{ij}e_{rs}]=[\frac{p}{n-1}]^{2}\binom{n-k}{2}k(n-k)\asymp p^{2}kn$
  5. 5.
    $\mathbb{E}[\sum_{i<j,r<s~{}:~{}x_{i}=x_{j}=1,x_{r}=x_{s}=-1}e_{ij}e_{rs}]=[\frac{p}{n-1}]^{2}\binom{k}{2}\binom{n-k}{2}\asymp p^{2}k^{2}$.
Plugging the above asymptotics in for the Binomials in Equation (52) suffices to obtain $\lVert H\rVert_{P}=\mathcal{O}(1)$.
Lemma B.4.
With high probability, the optimal objective value of (MaxCut-Hamming) satisfies:
$\mathcal{C}^{*}_{k}=\begin{cases}o(k\log(n))&\text{if}~{}k=o(n),\\ \mathcal{O}(n)&\text{if}~{}k=\Theta(n),\end{cases}$ (56)
where $k$ is the Hamming weight.
Proof.
We can use the indicator trick to try to bound the probability of a cut set of a given size. Let $z\in\{0,1\}^{n}$, $\lvert z\rvert=k$ and $I_{z,m}$ be a Bernoulli random variable indicating whether there are $m$ edges cut with assignment $z$ over the random choice of graph. Then
$X_{m}:=\sum_{z\in\{0,1\}^{n},\lvert z\rvert=k}I_{z,m}$
is the number of in-constraint cuts of size $m$. Note that $I_{z,m}$ are not independent. For any graph $G$ drawn from $\mathcal{G}(n,p/n)$, we have
$\mathbb{P}[I_{z,m}=1]=\binom{k(n-k)}{m}(p/n)^{m}(1-p/n)^{k(n-k)-m}.$
The first moment method gives:
$\mathbb{P}[X_{m}>0]\leq\mathbb{E}[X_{m}]\leq\binom{n}{k}\binom{k(n-k)}{m}n^{-m}.$
Suppose that $k=o(n)\cap\omega(1)$ and $m=\mathcal{O}(n)$. Then $k(n-k)=\mathcal{O}(nk)$, and we can apply the following asymptotics:
$\mathbb{P}[X_{m}>0]\leq\binom{n}{k}\binom{k(n-k)}{m}n^{-m}\asymp\left(\frac{n}{k}\right)^{k}\left(\frac{nk}{m}\right)^{m}n^{-m}=n^{k}k^{m-k}m^{-m}.$
The goal is to try to identify the phase transition point at which the probability goes to zero asymptotically. We can look at
$k\log(n)+m\log(k)-k\log(k)-m\log(m).$
Upon taking $m=\Theta(k\cdot r)$, we obtain
$k\log(n)+m\log(k)-k\log(k)-m\log(m)=k\log\left(\frac{n}{kr^{r}}\right).$
Transition is at $\log(n/k)=r\log(r)$. For $r=\log(n)$, $\mathbb{P}[X_{m}>0]\rightarrow 0$, thus $C^{*}_{k}=o(k\log(n))$.
Suppose $k=\Theta(n)$, then $\log\binom{n}{k}\asymp\mathcal{H}(\frac{k}{n})n$, where $\mathcal{H}$ is the binary entropy function. Thus
$\mathbb{P}[X_{m}>0]\leq\binom{n}{k}\binom{k(n-k)}{m}n^{-m}\asymp 2^{\mathcal{H}(n/k)n}\left(\frac{nk}{m}\right)^{m}n^{-m}=2^{\mathcal{H}(n/k)n}k^{m}m^{-m}.$
We can look at
$\mathcal{H}(n/k)n-m\log(k/m),$
so transition is at $m=2^{\mathcal{H}(n/k)}k=\Theta(n)$. Thus for $k=\Theta(n)$, $\mathcal{C}^{*}_{k}=\mathcal{O}(n)$. ∎
Appendix C Constrained Short Path via Penalized Objective
Suppose we want to solve the following constrained problem
$\min_{x\in\mathcal{X}\subseteq\{-1,1\}^{n}}H(x).$
Suppose we also have a CSP
$\mathcal{C}(x)=\sum_{\ell=1}^{m}C_{\ell}(x),$
with $m$ constraints $C_{\ell}$, indicating the in-constraint solutions. Suppose $C_{\ell}$ has $k_{\ell}$ literals and has $s_{\ell}$ satisfying assignments, then
$\mathcal{C}_{\ell}(x)=\begin{cases}-\frac{1}{s_{\ell}}\quad\mbox{if x satisfies constraint $\ell$}\\ \frac{1}{2^{k_{\ell}}-s_{\ell}}\quad\mbox{otherwise}.\end{cases}$
Consider the task of minimizing the penalized Hamiltonian
$\tilde{H}(x)=\frac{H(x)}{\lVert H\rVert_{2}}+\mathcal{C}(x).$ (57)
The Markov Chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is the random walk on the hypercube, and hence $\pi$ is the uniform distribution over $\{-1,1\}^{n}$. Thus, the short-path Hamiltonian is the same as [DPCB23] but with the penalized cost Hamiltonian.
The global minimum of Equation (57) is in fact the in-constraint minimum and $\mathbb{E}_{\pi}[\tilde{H}(x)]=\mathbb{E}_{\pi}[\frac{H(x)}{\lVert H\rVert_{2}}]$. Also note that $\lvert E^{*}\rvert=\Theta(m)$ for $\tilde{H}$. The below results will show that due to the normalization, the properties of the CSP are the only components that matter for determining the runtime. We denote $k=\max k_{\ell}$.
Lemma C.1.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the random walk on the Hamming hypercube. The penalized Hamiltonian is $\Delta_{P}$ stable with
$\Delta_{P}(\eta)=\mathcal{O}\left(\frac{mk}{n}\right).$ (58)
Proof.
$\Delta_{P}(\eta)\leq\frac{\mathbb{E}_{y\sim}[H(y)-H(x)]}{\lVert H\rVert_{2}}+\frac{mk}{n}=\mathcal{O}\left(\frac{mk}{n}\right).$ (59)
Lemma C.2 (Tail bound condition for Penalized Hamiltonian).
Suppose $\mathbb{E}_{\pi}[\tilde{H_{c}}]=0$,
$\pi((1-\eta)E^{\ast})\leq 2^{-\gamma},\quad\gamma=\Omega\left[\frac{2(1-\eta)^{2}}{\ell^{2}}\right].$
Proof.
Suppose flipping one variable $x_{i}$ changes at most $\Delta_{i}$ in the value of $H$ while holding other variables constant. Then
$\Delta_{i}=\mathcal{O}\left(\frac{mk}{n}\right).$
Assuming $\mathbb{E}_{\pi}[H(x)]=0$. Let $C(E)$ denote the number of $x$ with energy $\leq E$ under $\tilde{H}$. For $E<0$, $C(E)/2^{n}$ is the probability for a random $x$, $H(x)$ deviates below $\mathbb{E}_{\pi}[H(x)]=0$ by amount at least $|E|$. The $P$-pseudo Lipschitz norm of $\tilde{H}$ is bounded by
$\lVert\tilde{H}\rVert_{P}=\sum_{i=1}^{m}\frac{\Delta_{i}^{2}}{n}=\mathcal{O}\left(\frac{m^{2}k^{2}}{n^{2}}\right).$
Also, $\omega=\frac{2}{n}$ for the hypercube walk, and thus the Herbst argument implies
$\pi(E)\leq e^{-\Omega\left(\frac{2nE^{2}}{(mk)^{2}}\right)}.$
Taking $E=E^{\ast}(1-\eta)$, we have
$\pi((1-\eta)E^{\ast})\leq 2^{-n\gamma},\quad\gamma=\Omega\left[\left(\frac{{\left|E^{\ast}\right|}}{m}\right)^{2}\frac{2(1-\eta)^{2}}{k^{2}}\right].$
However $\lvert E^{\ast}\rvert=\Theta(m)$.
The following result is then evident from Theorem 4.3, given that from [DPCB23] we have that $b^{*}$ is constant for the random walk on the Hamming cube if $\gamma$ is constant.
Theorem C.3.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be random walk on the $n$-bit Hamming cube. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
$E^{\ast}:=\min_{x\in\mathcal{X}}H(x).$
If the number of constraints is $m=\Theta(n)$, then there exists a short-path algorithm with runtime
$\mathcal{O}^{\ast}\left(2^{n\left(\frac{1}{2}-\frac{(1-\eta)\lvert E^{\ast}\rvert b}{n2\ln(2)\Delta_{P}}\right)}\right).$
The penalty method ensures that our algorithm only outputs feasible solutions, but the framework only provides a speedup over unconstrained bruteforce search. Hence, the generalized short-path algorithm is significantly more effective.
Appendix D Details about Mixer Implementation
D.1 Block-encoding the Glauber Mixer
Since the Glauber dynamics mixer is a symmetric sparse matrix, we can implement the block-encoding in polynomial time by assuming sparse oracle access to the non-zero entries of $P$. Let $s$ be the maximum number of non-zero entries in any row of $D$. Then, we can implement the following oracle,
[…]
where $y$ is an index of a non-zero entry in $P(x,\cdot)$. Since Glauber dynamics can only update one site at most, the transition matrix contains at most $n$ entries. Hence, we can implement this oracle by computing $P$ at most $n$ times. Define the oracle for access to the elements of $P$ in the following way,
[…]
Implementing this oracle takes at most $\mathcal{O}(n)$. Finally, we define SWAP operator,
[…]
Then, the circuit $O_{S}^{\dagger}O_{A}^{\dagger}\operatorname{\textup{{SWAP}}}O_{A}O_{S}$ implements block-encoding of $D/s$. To see this, compute
[…]
One has
[…]
On the other hand,
[…]
Therefore,
[…]
Hence, we can implement the block-encoding in at most polynomial time.
D.2 Ground State Preparation for Glauber Mixer
We discuss how to prepare the ground state of discriminant operator for Glauber dynamics for hard-core model and Ising model at $\lambda\leq\lambda_{c}$ and $\beta_{c}\leq\beta$ respectively. This is all we can afford to prepare since the spectral gap of $D$ falls exponentially fast when $\lambda>\lambda_{c}$ ($\beta>\beta_{c}$) due to statistical phase transitions. For simplicity, we consider the hard-core model to explain the idea. However, the details for both models will be given as seperate prpositions below. Let $D_{\lambda}$ be the discriminant matrix of Glauber dynamics at fugacity $\lambda$. We use the block-encoding of $D_{\lambda}$ and amplitude amplification to prepare its ground state. Although, we can efficiently create block-encoding for $D_{\lambda}$ and apply singular value transformations to build a projector to its ground state, amplitude amplification might need exponentially many calls to this projector when we do not have a warm initial state. Instead, we combine classical annealing with quantum singular value transformation to prepare the ground state of $D_{\lambda}$ to create sequence of states where each state is warm with respect to the next one. Suppose that we prepare $\pi^{(1)}$, the coherent quantum state corresponding to the ground state of $D_{\lambda_{1}}$ where $0<\lambda_{1}<\varepsilon$,
|$\ket{\pi^{(1)}}=\sum_{x\in\mathcal{X}}\sqrt{\pi_{\lambda_{1}}(x)}\ket{x}$
where the sum is over all independent sets $x$ in $G$. Next, we increase fugacity to $\lambda_{2}=\lambda_{1}(1+\Delta)$ and prepare,
|$\ket{\pi^{(2)}}=\sum_{x\in\mathcal{X}}\sqrt{\pi_{\lambda_{2}}(x)}\ket{x}.$
This quantum state can be prepared by applying ground state projector of $D_{\lambda_{1}}$ to and $D_{\lambda_{2}}$ to |$\ket{\pi^{(0)}}$ through fixed point amplitude amplification. We repeat this process until we prepare |$\ket{\pi^{(k)}}$,
|$\ket{\pi^{(k)}}=\sum_{x\in\mathcal{X}}\sqrt{\pi_{\lambda_{k}}(x)}\ket{x}$
with $\lambda_{k}=\lambda_{c}$. We need to show that this process can be done in $\operatorname{poly}(n)$ time.
Proposition D.1 (Preparation of Gibbs State for Hard-core Model).
Consider Glauber dynamics chain for hard-core model on graph $G(\mathcal{N},\mathcal{E})$ with transition matrix $P_{\lambda}$ with stationary distribution,
$\pi_{\lambda}(x)=\frac{\lambda^{|x|}}{Z}.$ (60)
Let $\delta(\lambda)>0$ be the spectral gap of the discriminant matrix $D(P_{\lambda})$ associated with $P_{\lambda}$. Then, there exists a quantum algorithm that prepares the ground state of $-D$ up to $\varepsilon$ accuracy in $L2$ norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\lambda^{\prime}\leq\lambda}\delta(\lambda^{\prime})$.
Proof.
We start with the following quantum state,
|$\ket{\pi^{(0)}}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ket{x_{i}}$
where $x_{i}$ is all 00 bit string except location $i$. This quantum state is ground state of $D$ at $\lambda=0$ as each $x_{i}$ is an independent set and they are the only ones with non-zero probability. This quantum state is essentially superposition over all strings with Hemming weight 1. This state can be prepared by applying $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}$ where $X_{i}$ is Pauli-$X$ applied to the all 00 state. Therefore, this state can be prepared in polynomial time. However, we cannot implement the Glauber dynamics at $\lambda=0$ as the transition density is not meaningful. Instead, we start with |$\ket{\pi^{(1)}}$ which is the ground state of $D_{\lambda_{1}}$ which can be prepared from |$\ket{\pi^{(0)}}$ since |$\ket{\pi^{(0)}}$ and |$\ket{\pi^{(1)}}$ overlaps significantly
$|\braket{\pi^{(0)}}{\pi^{(1)}}|=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{\lambda_{1}^{1/2}}{\sqrt{Z_{1}}}\geq\frac{\sqrt{n\lambda_{1}}}{\sqrt{n\lambda_{1}+\sum\limits_{|\mathcal{I}|>1}\lambda_{1}^{\mathcal{|I|}}}}$ $=\Omega(1),$
for $\lambda_{1}\leq\frac{3n}{2^{n}}$. Similarly, given |$\ket{\pi^{(k-1)}}$, we can prepare |$\ket{\pi^{(k)}}$ efficiently. The number of calls to the ground state projector by the fixed point amplitude amplification from |$\ket{\pi^{(k-1)}}$ to |$\ket{\pi^{(k)}}$ is $\widetilde{\mathcal{O}}(|\langle\pi^{(k-1)}|\pi^{(k)}\rangle|^{-1})$. The overlap can be calculated as
$\left|\!\braket{\pi^{(k-1)}}{\pi^{(k)}}\!\right|$ | $=\sum_{x\in\mathcal{X}}\frac{\lambda_{k-1}^{|x|/2}}{\sqrt{Z_{k-1}}}\frac{\lambda_{k}^{|x|/2}}{\sqrt{Z_{k}}}$
$\geq\sum_{x\in\mathcal{X}}\frac{\lambda_{k-1}^{|x|/2}\lambda_{k}^{|x|/2}}{Z_{k}}$
$=\sum_{x\in\mathcal{X}}\frac{\lambda_{k}^{|x|}\Delta^{-|x|/2}}{Z_{k}}$
$\geq(1+\Delta)^{\frac{-n}{2}}$
$\geq 1-\frac{n\Delta}{2}$
$=\Omega(1),$
if we set $\Delta\leq\frac{2}{n}$. Hence, starting from |$\ket{\pi^{(1)}}$ we can prepare a schedule that maintains constant overlap with the subsequent state. Since we have constant overlap throughout the schedule, each amplitude amplification step only requires constant number of calls to ground state projectors. Also note that implementing the ground state projector requires $\widetilde{\mathcal{O}}(\delta^{-1})$ calls to block-encoding of $D$ [GSLW19]. Finally, we only need to do $\mathcal{O}(\operatorname{poly}(n))$ rounds of amplitude amplification since $\lambda_{k}=\lambda_{1}\exp(2k/n)$ and for $k\geq\frac{n}{2}\log(\lambda_{c}/\lambda_{1})$, $\lambda_{k}\geq\lambda_{c}$.
Proposition D.2 (Preparation of Gibbs State for Ising Model Model).
Consider Glauber dynamics chain for Ising model on graph $G(\mathcal{N},\mathcal{E})$ with transition matrix $P_{\beta}$ with stationary distribution,
$\pi_{\beta}(x)=\frac{\exp(-\beta H(x))}{Z}.$ (61)
Let $\delta(\beta)>0$ be the spectral gap of the discriminant matrix $D(P_{\beta})$ associated with $P_{\beta}$. Then, there exists a quantum algorithm that prepares the ground state of $-D$ up to $\varepsilon$ accuracy in $\ell_{2}$-norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\beta^{\prime}\leq\beta}\delta(\beta^{\prime})$.
Proof.
The proof is similar to the hard-core model, and we start with the following quantum state,
|$\ket{\pi^{(0)}}=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}\ket{x}$
This quantum state is ground state of $D$ at $\beta=0$ since for $\beta=0$, Glauber dynamics is equivalent to hypercube walk. Similar to MIS case, given |$\ket{\pi^{(k-1)}}$, we can prepare |$\ket{\pi^{(k)}}$ up to $\varepsilon$ accuracy efficiently. The number of calls to the ground state projector by the fixed point amplitude amplification from |$\ket{\pi^{(k-1)}}$ to |$\ket{\pi^{(k)}}$ is $\widetilde{\mathcal{O}}(|\langle\pi^{(k-1)}|\pi^{(k)}\rangle|^{-1})$. The overlap can be calculated as
$\left|\!\braket{\pi^{(k-1)}}{\pi^{(k)}}\!\right|$ | $=\sum_{x\in G}\frac{\exp(-\beta_{k-1}H(x)/2)}{\sqrt{Z_{k-1}}}\frac{\exp(-\beta_{k}H(x)/2)}{\sqrt{Z_{k}}}$
$\geq\sum_{x\in G}\frac{\exp(-\beta_{k-1}H(x)/2)\exp(-\beta_{k}H(x)/2)}{Z_{k}}$
$=\sum_{x\in G}\frac{\exp(-\beta_{k}H(x))\exp((\beta_{k}-\beta_{k-1})H(x)/2)}{Z_{k}}$
$\geq\exp((\beta_{k}-\beta_{k-1})\|H\|/2)$
$=\Omega(1),$
if we set $(\beta_{k}-\beta_{k-1})=\frac{1}{\|H\|}$. Hence, starting from |$\ket{\pi^{(0)}}$ we can prepare a schedule that maintains constant overlap with the subsequent state.
Since we have constant overlap throughout the schedule, each amplitude amplification step only requires constant number of calls to ground state projectors. Also note that implementing the ground state projector requires $\widetilde{\mathcal{O}}(\delta^{-1})$ calls to block-encoding of $D$ [GSLW19]. Finally, we only need to do $\operatorname{poly}(n)$ rounds of amplitude amplification since $\|H\|=\operatorname{poly}(n)$ and length of the schedule is polynomial. ∎